### v2

# Simplifying Hypergraph Neural Networks

Abstract
Hypergraphs, with hyperedges connecting multiple nodes, are crucial for modelling higher-order interactions in real-world data. In frameworks utilising hypergraphs for downstream tasks, a task-specific model is typically paired with a hypergraph neural network (HNN). HNNs enhance the task-specific model by generating node features with hypergraph structural information via message passing. However, the training for HNNs is often computationally intensive, which limits their practical use. To tackle this challenge, we propose an alternative approach by integrating hypergraph structural information into node features using a training-free model called simplified hypergraph neural network (SHNN) that only contains a predefined propagation step. We theoretically show the efficiency and effectiveness of SHNN by showing that: 1) It largely reduces the training complexity when solving hypergraph-related downstream tasks compared to existing HNNs; 2) It utilises as much information as existing HNNs for node feature generation; and 3) It is robust against the oversmoothing issue while using long-range interactions. Experiments in node classification and hyperedge prediction showcase that, compared to state-of-the-art HNNs, SHNN leads to both competitive performance and superior training efficiency. Notably, on Cora-CA, the SHNN-based framework achieves the highest node classification accuracy with just $2\%$ training time of the best baseline.
Machine Learning, ICML
1 Introduction
Higher-order interactions involving more than two entities exist in various domains, such as co-authorships in social science (Han et al., 2009) and the spreading phenomena in epidemiology (Jhun, 2021). To model these complicated interactions, hypergraphs, an extension of traditional graphs, are widely employed (Bick et al., 2023). Hypergraphs consist of nodes representing entities and hyperedges denoting higher-order interactions that connect multiple entities. Recently, hypergraph neural networks (HNNs) have leveraged message passing over hypergraph structures to enhance node feature learning, achieving remarkable performance in various downstream tasks like node classification (Feng et al., 2019; Bai et al., 2021; Chien et al., 2022; Wang et al., 2023a; Huang & Yang, 2021; Wang et al., 2023b; Duta et al., 2023; Telyatnikov et al., 2023). However, the training of HNNs is typically computationally intensive, which limits their practical applications, especially in semi-supervised learning that utilises both training and test data simultaneously (Joachims et al., 1999; Chapelle et al., 2009).
To address this challenge, we take inspiration from the literature on scalable and efficient graph neural networks (GNNs). In this literature, SIGN (Frasca et al., 2020) proposes to move the integration of graph information from the training stage to the data preprocessing phase by a training-free model, and SGC (Wu et al., 2019) builds an efficient architecture by collapsing multiple layers of graph convolutions with the intermediate non-linear activation functions removed. We aim to explore similar ideas for hypergraphs.
In this paper, we propose a novel model called simplified hypergraph neural network (SHNN), by simplifying four representative HNNs in two steps. First, we make the selected HNNs training-free by eliminating their learnable parameters. Next, we enhance inference efficiency by removing non-linear activation functions and converting an $L$-layer HNN into a single propagation step. Remarkably, this two-step simplification unifies the four chosen HNNs into a single formulation, despite their different underlying design philosophies. We refer to this formulation as SHNN. We further conduct a thorough analysis of SHNN to theoretically demonstrate its efficiency and effectiveness. Generally, we present three key properties of SHNN: 1) SHNN significantly reduces training complexity in hypergraph-related downstream tasks compared to existing HNNs; See Figure 1 for the comparison between the training pipelines of an SHNN and an HNN in node classification; 2) SHNN can utilise as much information as existing HNNs for generating node features; and 3) SHNN is robust against the oversmoothness issue when taking into account the long-range informations.
The contributions of this work are summarised as follows:
$\bullet$ We develop SHNN, an efficient and effective training-free model for generating node features with the structural information in pre-defined hypergraphs.
$\bullet$ We theoretically support the effectiveness and efficiency of SHNN by showing that: 1) It leads to remarkably low training complexity when solving hypergraph-related downstream tasks; 2) It utilises the same amount of information as existing HNNs for generating node features; and 3) It is robust against the oversmoothing issue when utilising long-range interactions.
$\bullet$ We conduct extensive experiments in both node classification and hyperedge prediciton tasks to compare SHNN with nine state-of-the-art HNNs. The empirical results show that the SHNN-based framework exhibits both the competitive performance and superior training efficiency.
Figure 1: Training pipeline of an HNN vs. an SHNN in node classification. Top row: The HNN uses a hypergraph structure to generate node features by $(L+1)$ learnable forward propagation blocks (FP Blocks), which are then used by a classifier, with the HNN and the classifier being trained simultaneously. Bottom row: The SHNN-based framework consists solely of the classifier being trained for node classification and the SHNN, which can be precomputed prior to classifier training.
2 Notation
Hypergraphs. Let =$\mathcal{H}=\{\mathcal{V},\mathcal{E},\mathbf{H}\}$ be a hypergraph, where $\mathcal{V}=\{v_{1},v_{2},\cdots,v_{n}\}$ is the node set, =$\mathcal{E}=\{e_{1},e_{2},\cdots,e_{m}\}$ is the hyperedge set, and $\mathbf{H}\in\{0,1\}^{n\times m}$ is an incidence matrix in which $\mathbf{H}_{ik}=1$ means that $e_{j}$ contains node $v_{i}$ and $\mathbf{H}_{ik}=0$ otherwise. Define $\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}\in\mathbb{R}_{\geq 0}^{n\times n}$ as a diagonal matrix of node degrees and $\mathbf{D}_{\mathcal{H}^{\mathcal{E}}}\in\mathbb{R}_{\geq 0}^{m\times m}$ as a diagonal matrix of hyperedge degrees, where $\mathbf{D}_{\mathcal{H}^{\mathcal{V}}_{ii}}$ and $\mathbf{D}_{\mathcal{H}^{\mathcal{E}}_{kk}}$ are the number of hyperedges with $v_{i}$ and the number of nodes in $e_{k}$, respectively. In the following discussions, we assume the hypergraph does not have isolated nodes or empty/duplicate hyperedges. Moreover, we define: 1) The distance between two nodes on a hypergraph is the number of hyperedges in the shortest path between them, e.g., the distance is one if they are directly connected by a hyperedge; 2) The $k$-hop neighbours of a node $v_{i}$ are all nodes with a distance of $k$ or less; and 3) The neighbouring nodes of a node are its 1-hop neighbours.
Graphs. Let $\mathcal{G}=\{\mathcal{V},\mathbf{W}\}$ be a graph, where $\mathcal{V}=\{v_{1},v_{2},\cdots,v_{n}\}$ is the node set, and $\mathbf{W}\in\mathbb{R}_{\geq 0}^{n\times n}$ is the adjacency matrix of $\mathcal{G}$ in which $\mathbf{W}_{ij}>0$ means that $v_{i}$ and $v_{j}$ are connected and $\mathbf{W}_{ij}=0$ otherwise. We set $\mathbf{D}\in\mathbb{R}_{\geq 0}^{n\times n}$ as a diagonal node degree matrix for $\mathcal{G}$, where $\mathbf{D}_{ii}$ is the sum of the $i$-th row of $\mathbf{W}$. Moreover, we denote the graph Laplacian of $\mathcal{G}$ as $\mathbf{L}=\mathbf{D}-\mathbf{W}$.
Clique expansion. Given a hypergraph =$\mathcal{H}=\{\mathcal{V},\mathcal{E},\mathbf{H}\}$, its clique expansion is defined as a graph $\mathcal{G}=\{\mathcal{V},\mathbf{W}\}$, where $\mathcal{V}$ remains unchanged, and $\mathbf{W}_{ij}>0$ if and only if $v_{i}$ and $v_{j}$ are connected by a hyperedge on $\mathcal{H}$ and $\mathbf{W}_{ij}=0$ otherwise. Hence, each hyperedge in $\mathcal{H}$ is a clique in $\mathcal{G}$.
Other notations. We set node features as $\mathbf{X}_{\mathcal{V}}\!=\![\mathbf{x}^{\top}_{v_{1}},\!\mathbf{x}^{\top}_{v_{2}},\!\cdots,\!\mathbf{x}^{\top}_{v_{n}}]^{\top}\in\mathbb{R}^{n\times d}$, which is a matrix that contains $d$-dimensional features. We denote functions or variables at the $l$-th layer of a model using the superscript $(l)$ and use $\oplus$ for concatenation. We use $\mathbf{\Theta}$ to represent the learnable weight matrix in a model, $\operatorname{MLP}(\cdot)$ for a multilayer perceptron (MLP), and $\sigma(\cdot)$ for non-linear functions (e.g., ReLU). In subsequent sections, we presume that the node features are $d$-dimensional unless otherwise noted.
3 Methodology
In this section, we first revisit existing hypergraph neural networks (HNNs), summarising these models along with their corresponding learning frameworks for downstream tasks into general formulations. Furthermore, we develop the simplified hypergraph neural network (SHNN) by removing the learnable parameters and the non-linear function from four representative HNNs.
3.1 Revisiting Hypergraph Neural Networks
Table 1: Overview for four state-of-the-art HNNs. In this table, $\gamma_{U},\gamma_{E},\gamma_{D},\gamma^{\prime}_{l}\!\in\!(0,1)$ are hyperparameters and $\mathbf{I}\in\mathbb{R}^{d\times d}$ denotes an identity matrix.
Name Type Message Passing Formulation
UniGCNII (Huang & Yang, 2021) Direct Message Passing $\mathbf{X}_{\mathcal{V}}^{(l)}\!=\!\sigma\Bigg{(}\Big{(}(\!1\!-\!\gamma_{U}\!)\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}^{-1/2}\mathbf{H}\!\tilde{\mathbf{D}}_{\mathcal{H}^{\mathcal{E}}}^{-1/2}\mathbf{D}_{\mathcal{H}^{\mathcal{E}}}^{-1}\mathbf{H}^{\top}\mathbf{X}_{\mathcal{V}}^{(l-1)}\!+\!\gamma_{U}\mathbf{X}_{\mathcal{V}}^{(0)}\Big{)}\mathbf{\Theta}^{(l)}\Bigg{)}$
Deep-HGNN (Chen et al., 2022) Direct Message Passing $\mathbf{X}_{\mathcal{V}}^{(l)}\!=\!\sigma\Bigg{(}\Big{(}(\!1\!-\!\gamma_{D}\!)\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}^{-1/2}\mathbf{H}\!\mathbf{D}_{\mathcal{H}^{\mathcal{E}}}^{-1}\mathbf{H}^{\top}\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}^{-1/2}\mathbf{X}_{\mathcal{V}}^{(l)}\!+\!\gamma_{D}\mathbf{X}_{\mathcal{V}}^{(0)}\Big{)}\Big{(}(1-\gamma^{\prime}_{D})\mathbf{I}+\gamma^{\prime}_{D}\mathbf{\Theta}^{(l)}\Big{)}\Bigg{)}$
AllDeepSets (Chien et al., 2022) Indirect Message Passing $\mathbf{X}_{\mathcal{V}}^{(l)}=\operatorname{MLP}\Bigg{(}\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}^{-1}\mathbf{H}\operatorname{MLP}\Big{(}\mathbf{D}_{\mathcal{H}^{\mathcal{E}}}^{-1}\mathbf{H}^{\top}\operatorname{MLP}(\mathbf{X}_{\mathcal{V}}^{(l-1)})\Big{)}\Bigg{)}$
ED-HNN (Wang et al., 2023a) Indirect Message Passing $\mathbf{X}_{\mathcal{V}}^{(l)}=\operatorname{MLP}\Bigg{[}(1-\gamma_{E})\operatorname{MLP}\Bigg{(}\mathbf{D}^{-1}_{\mathcal{H}^{\mathcal{V}}}\mathbf{H}\operatorname{MLP}\Big{(}\mathbf{D}^{-1}_{\mathcal{H}^{\mathcal{E}}}\mathbf{H}^{\top}\operatorname{MLP}(\mathbf{X}_{\mathcal{V}}^{(l-1)})\Big{)}\Bigg{)}+\gamma_{E}\mathbf{X}_{\mathcal{V}}^{(0)}\Bigg{]}$
Model design. Generally, existing HNNs are based on the message passing mechanism, adapted from the graph machine learning literature (Xu et al., 2019; Bronstein et al., 2021). This mechanism makes models iteratively update the features of each node by aggregating the features of its neighbours. Mathematically, the features of a specific node $v_{i}$ are updated by the following propagation operators:
$\mathbf{x}_{v_{i}}^{(0)}=\operatorname{MLP}(\mathbf{x}_{v_{i}}),~{}~{}\mathbf{x}_{v_{i}}^{(l)}=\phi_{\mathbf{\Theta}}(\mathbf{x}_{v_{i}}^{(0)},\mathbf{x}_{v_{i}}^{(l-1)},\oplus_{v_{j}\in\mathcal{N}_{\mathcal{H}_{v_{i}}}}\mathbf{x}_{v_{j}}^{(l-1)}),$ (1)
where $\phi_{\mathbf{\Theta}}(\cdot)$ denotes a learnable non-linear function that facilitates the message passing between neighbouring nodes, and $\mathcal{N}_{\mathcal{H}_{v_{i}}}$ is the set of $1$-hop neighbours of $v_{i}$ on the hypergraph. Specifically, $\phi_{\mathbf{\Theta}}(\cdot)$ can be categorized into two types: direct and indirect message passing functions. The direct approaches aggregate features of neighbouring nodes to that of the target node directly (Huang & Yang, 2021; Wang et al., 2023b; Chen et al., 2022). In the indirect approaches, the features of neighbouring nodes are aggregated to that of the target node via some virtual nodes (Chien et al., 2022; Wang et al., 2023a; Yang et al., 2022). Mathematically these approaches can be formulated as:
$\mathbf{x}_{v_{i}}^{(l)}$ $=f_{3}\Big{(}f_{0}(\mathbf{x}_{v_{i}}^{(0)})+f_{1}(\mathbf{x}_{v_{i}}^{(l-1)})+\sum_{v_{j}\in\mathcal{N}_{\mathcal{G}_{v_{i}}}}f_{2}(\mathbf{x}^{(l-1)}_{v_{j}})\Big{)},$ (2a)
$\mathbf{x}_{v_{i}}^{(l)}$ $=g_{2}\Big{(}g_{0}(\mathbf{x}_{v_{i}}^{(0)})+\mathbf{p}^{\top}g_{1}(\mathbf{x}_{v_{i}}^{(l-1)},\oplus_{v_{j}\in\mathcal{N}_{\mathcal{H}_{v_{i}}}}\mathbf{x}_{v_{j}}^{(l-1)})\Big{)},$ (2b)
where $f_{0}(\cdot)$, $f_{1}(\cdot)$, $f_{2}(\cdot)$, $f_{3}(\cdot)$, $g_{0}(\cdot)$, $g_{1}(\cdot)$, $g_{2}(\cdot)$ represent seven learnable non-linear functions, and $\mathbf{p}\in\mathbb{R}^{n_{v}}$ is a learnable vector used to aggregate the features of virtual nodes generated by $g_{1}(\cdot)$. Here, $n_{v}$ is a hyperparameter corresponding to the number of the virtual nodes, and Eq. (2a) and Eq. (2b) represent the direct and indirect approaches respectively. Further, we detail four representative HNNs in Table 1, and show how specific HNNs are special cases of Eq. (1) in Appendix.
Application for downstream tasks. The pipeline of an HNN-based learning framework for downstream tasks can be formulated as follows:
$\hat{\mathbf{Y}}=\varphi_{\mathbf{\Theta}}(\hat{\mathbf{X}}_{\mathcal{V}}),~{}~{}\hat{\mathbf{X}}_{\mathcal{V}}=\Phi_{\mathbf{\Theta}}(\mathbf{X}_{\mathcal{V}},\mathbf{H}),$ (3)
where $\varphi_{\mathbf{\Theta}}(\cdot)$ is a task-specific model, $\hat{\mathbf{Y}}$ denotes the task-specific output whose dimension is task-dependent [^1] For instance, in the node classification task, $\hat{\mathbf{Y}}\in\mathbb{R}^{n\times c}$ contains the logits for $c$ categories; and in the hyperlink prediction task, $\hat{\mathbf{Y}}\in\mathbb{R}^{m_{p}}$ contains the probability of $m_{p}$ potential hyperedges., $\hat{\mathbf{X}}_{\mathcal{V}}\in\mathbb{R}^{n\times d}$ denote the features generated by an HNN, and $\Phi_{\mathbf{\Theta}}(\cdot)$ is an HNN that updates the node features with the forward propagation functions formulated as Eq. (1). The models $\varphi_{\mathbf{\Theta}}(\cdot)$ and $\Phi_{\mathbf{\Theta}}(\cdot)$ are trained together with the supervision given by the downstream tasks. As illustrated in Table 1, existing HNNs contain complex learnable message passing formulation, which makes the training of the HNN-based framework computationally expensive.
3.2 Simplifying Hypergraph Neural Networks
In this section, our goal is to introduce the simplified hypergraph neural network (SHNN) by simplifying the representative HNNs presented in Table 1.
Learnable parameters elimination. To develop a training-free model, we first eliminate the learnable parameters from the message passing formulation of four selected HNNs. Specifically, we replace the learnable matrices in their formulations with identity matrices. This simplification allows us to represent the simplified formulations of the chosen HNNs as follows:
$\mathbf{X}_{\mathcal{V}}^{(l)}$ $=\sigma\Big{(}(1-\gamma_{U})\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}^{-1/2}\mathbf{H}\!\tilde{\mathbf{D}}_{\mathcal{H}^{\mathcal{E}}}^{-1/2}\mathbf{D}_{\mathcal{H}^{\mathcal{E}}}^{-1}\mathbf{H}^{\top}\mathbf{X}_{\mathcal{V}}^{(l-1)}+\gamma_{U}\mathbf{X}_{\mathcal{V}}^{(0)}\Big{)},$ (4a)
$\mathbf{X}_{\mathcal{V}}^{(l)}$ $=\sigma\Big{(}(1-\!\gamma_{D}\!)\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}^{-1/2}\mathbf{H}\!\mathbf{D}_{\mathcal{H}^{\mathcal{E}}}^{-1}\mathbf{H}^{\top}\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}^{-1/2}\mathbf{X}_{\mathcal{V}}^{(l-1)}\!+\!\gamma_{D}\mathbf{X}_{\mathcal{V}}^{(0)}\Big{)},$ (4b)
$\mathbf{X}_{\mathcal{V}}^{(l)}$ $=\sigma\Bigg{(}\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}^{-1}\mathbf{H}\sigma\Big{(}\mathbf{D}_{\mathcal{H}^{\mathcal{E}}}^{-1}\mathbf{H}^{\top}\sigma(\mathbf{X}_{\mathcal{V}}^{(l-1)})\Big{)}\Bigg{)},$ (4c)
$\mathbf{X}_{\mathcal{V}}^{(l)}$ $=\sigma\Bigg{[}(1-\gamma_{E})\sigma\Bigg{(}\mathbf{D}^{-1}_{\mathcal{H}^{\mathcal{V}}}\mathbf{H}\sigma\Big{(}\mathbf{D}^{-1}_{\mathcal{H}^{\mathcal{E}}}\mathbf{H}^{\top}\sigma(\mathbf{X}_{\mathcal{V}}^{(l-1)})\Big{)}\Bigg{)}+\gamma_{E}\mathbf{X}_{\mathcal{V}}^{(0)}\Bigg{]},$ (4d)
where Eq. (4a), Eq. (4b), Eq. (4c) and Eq. (4d) represent the simplified formulation of UniGCNII, Deep-HGNN, AllDeepSets and ED-HNN, respectively. Notably, $\sigma(\cdot)$ only denotes a general non-linear function without parameters, not implying that these simplified models use the same function.
Linearisation. Recently, some works in graph machine learning (Wu et al., 2019; Wang & Zhang, 2022) show that the message passing mechanism can operate more efficiently without non-linear activation functions while maintaining high effectiveness. Inspired by these works, we further remove the non-linearity from the formulations represented by Eqs. (4$\sim$4d), thereby transforming an L-layer HNN into a single propagation step. For the sake of brevity, we present the simplification results as Lemma 3.1 and provide the mathematical proof for this Lemma in Appendix.
Lemma 3.1.
Let $\alpha\in[0,1)$ be a hyperparameter, $\mathbf{X}_{\mathcal{V}}$ be the input node features, and $\hat{\mathbf{X}}_{\mathcal{V}}$ be the output of an $L$-layer UniGCNII/AllDeepSets/ED-HNN/Deep-HGNN. Assume that the learnable parameters and the non-linearity are removed from the model. Then, for any hypergraph =$\mathcal{H}=\{\mathcal{V},\mathcal{E},\mathbf{H}\}$, there exists a clique expansion $\mathcal{G}=\{\mathcal{V},\mathbf{W}\}$ to make $\hat{\mathbf{X}}_{\mathcal{V}}=\Big{(}(1-\alpha)^{L}\mathbf{W}^{L}+\alpha\sum_{l=0}^{L-1}(1-\alpha)^{l}\mathbf{W}^{l}\Big{)}\mathbf{X}_{\mathcal{V}}$.
According to Lemma 3.1, we design the formulation of our SHNN as the following equation:
$\hat{\mathbf{X}}=\mathbf{S}\mathbf{X}_{\mathcal{V}},$ (5)
where $\mathbf{S}=(1-\alpha)^{L}\mathbf{W}^{L}+\alpha\sum_{l=0}^{L-1}(1-\alpha)^{l}\mathbf{W}^{l}$ and we define that the $L$ used to compute $\mathbf{S}$ is the number of layers of an SHNN. To intuitively understand the behaviour of an $L$-layer SHNN, we present the Lemma 3.2. For conciseness, we show the proof of this Lemma in Appendix.
Lemma 3.2.
Let $\mathcal{H}=(\mathcal{V},\mathcal{E},\mathbf{H})$ be a hypergraph and $\mathcal{G}=(\mathcal{V},\mathbf{W})$ be its clique expansion with self-loops. Assume that $\alpha\in[0,1)$, and $\mathbf{S}\!=\!(1-\alpha)^{L}\mathbf{W}^{L}+\alpha\sum_{l=0}^{L-1}(1-\alpha)^{l}\mathbf{W}^{l}$. Then, we have $\mathbf{S}_{ij}>0$ if and only if $v_{i}$ is a $L$-hop neighbour of $v_{j}$ on $\mathcal{H}$, and $\mathbf{S}_{ij}=0$ otherwise.
By Lemma 3.2, an $L$-layer SHNN enables the information exchange between any node $v_{i}$ and its $L$-hop neighbours on the hypergraph. In the next paragraph, we detail the specific designs of $\mathbf{S}$.
Operator design. To apply the SHNN to hypergraph-related downstream tasks, the key is designing $\mathbf{W}$ to generate $\mathbf{S}$. We achieve this using a hyperedge-size-based edge weight that is defined as:
$\mathbf{W}_{{H}_{ij}}=\sum_{k=1}^{m}\frac{\delta(v_{i},v_{j},e_{k})}{\mathbf{D}_{\mathcal{H}^{\mathcal{E}}_{kk}}},$ (6)
where $\delta(\cdot)$ is a function that returns 1 if $e_{k}$ connect $v_{i}$ and $v_{j}$ and returns 0 otherwise. The main intuition behind the design of this edge weight is that, under certain conditions, the value of $\mathbf{W}_{H_{ij}}$ is positively correlated with the probability of nodes $v_{i}$ and $v_{j}$ having the same label. Consequently, this edge weight design can make the weighted clique expansion fit the homophily assumption, which is that connected nodes tend to be similar to each other (McPherson et al., 2001). We theoretically elaborate on this intuition in Appendix. Furthermore, inspired by previous works in the graph domain (Wu et al., 2020), we generate $\mathbf{S}$ based on a symmetrically normalised $\mathbf{W}_{H}$ with the self-loop. Specifically, we generate the $\mathbf{S}$ by:
$\mathbf{S}=(1-\alpha)^{L}(\tilde{\mathbf{D}}_{H}^{-1/2}\tilde{\mathbf{W}}_{H}\tilde{\mathbf{D}}_{H}^{-1/2})^{L}+\alpha\sum_{l=0}^{L-1}(1-\alpha)^{l}(\tilde{\mathbf{D}}_{H}^{-1/2}\tilde{\mathbf{W}}_{H}\tilde{\mathbf{D}}_{H}^{-1/2})^{l},$ (7)
where $\tilde{\mathbf{W}}_{H}=\mathbf{W}_{H}+\mathbf{I}_{n}$ and $\tilde{\mathbf{D}}_{H}\in\mathbb{R}^{n\times n}$ is the diagonal node degree matrix of $\tilde{\mathbf{W}}_{H}$. In the next section, we present the theoretical analysis to support the efficiency and effectiveness of our SHNN.
4 Analysis
In this section, we begin by showing the efficiency of the proposed SHNN through a comparison of the training complexity of the SHNN-based framework with the HNN-based framework in hypergraph-related downstream tasks. Additionally, we demonstrate the effectiveness of our SHNN by analyzing its information utilisation and its robustness against the oversmoothing issue.
Training complexity in downstream tasks. According to Eq. (5), the SHNN-based framework for downstream tasks can be summarised as follows:
$\hat{\mathbf{Y}}=\varphi_{\mathbf{\Theta}}(\mathbf{S}\mathbf{X}_{\mathcal{V}}).$ (8)
As illustrated by Eq. (8), the SHNN-based framework for downstream tasks requires training only the task-specific model. Additionally, since the SHNN relies on a pre-defined one-step propagation operator, its computations can be handled during data pre-processing. Consequently, the training complexity of the SHNN-based framework is determined solely by the task-specific model. In contrast, on the basis of Eq. (3), the HNN-based framework involves training both the HNN and the task-specific model. See Figure 1 for an example that compares the training pipelines of the SHNN-based and HNN-based framework. The HNN includes complex learnable operations that facilitate message passing between connected nodes on a hypergraph, inherently making the training of the HNN-based framework more complex than the SHNN-based framework. To provide a precise comparison, we summarize the training complexity of the SHNN-based framework and the four frameworks based on different state-of-the-art HNNs in Table 2. This table shows that the training complexity of the SHNN-based framework is consistently lower than that of the HNN-based frameworks, theoretically confirming the training efficiency of the SHNN-based framework.
Table 2: Training computational complexity of SHNN/HNN-based frameworks. Here, $M$ denote the training computational complexity of the task-specific model, $n$ is the number of nodes, $m$ is the number of hyperedges, $m^{\prime}$ is the number of edges in the clique expansion, $\|\mathbf{H}\|_{0}$ is the number of non-zero entities in $\mathbf{H}$, $L$ is the number of layers, and $d$ is the feature dimension.
SHNN UniGCNII AllDeepSets Deep-HGNN ED-HNN
$\mathcal{O}(M)$ $\mathcal{O}(M+TL(n+m+\|\mathbf{H}\|_{0})d+TLnd^{2})$ $\mathcal{O}(M+TL\|\mathbf{H}\|_{0}d+TL(n+m)d^{2})$ $\mathcal{O}(M+TLm^{\prime}d+TLnd^{2})$ $\mathcal{O}(M+TL\|\mathbf{H}\|_{0}d+TL(n+m)d^{2})$
Information utilisation. The core strength of existing HNNs lies in their ability to utilize the information embedded within hypergraph structures to generate powerful node features for downstream tasks. Consequently, we analyse the effectiveness of our proposed SHNN by comparing its information utilization capabilities with those of existing HNNs, which is summarised in Proposition 4.1.
Proposition 4.1.
Let $H_{0}$ denote the entropy of the information that an HNN utilizes by $L$ message passing function $\phi_{\mathbf{\Theta}}(\cdot)$ introduced in Eq. (1) for generating node features. Similarly, let $H_{1}$ denote the entropy of the information that an $L$-layer SHNN uses for the same purpose. Then, $H_{0}=H_{1}$.
Proof Sketch.
To prove this proposition, we first let $H_{0}^{v_{i}}$ be the entropy of information used by an HNN with $L$ message passing layers in generating the features of node $v_{i}$, $H_{1}^{v_{i}}$ be the entropy of information used by an $L$-layer SHNN for the same purpose, and $H_{2}^{v_{i}}$ denote the entropy of information within node $v_{i}$ and its $L$-hop neighbours. Theron, we prove this proposition by showing that $H_{0}^{v_{i}}=H_{2}^{v_{i}}$ and $H_{1}^{v_{i}}=H_{2}^{v_{i}}$. The detailed proof is presented in Appendix. ∎
This proposition theoretically demonstrates that SHNN is as effective as existing HNNs in utilizing information from pre-defined hypergraphs to generate node features. Specifically, both SHNN and existing HNNs achieve this by aggregating neighbourhood information.
Robustness against the oversmoothing issue. According to Lemma 3.2, SHNN can leverage global interactions within the given hypergraph to generate node features by increasing the number of model layers. However, as highlighted in a recent study (Chen et al., 2022), message-passing-based models on hypergraphs may suffer from the oversmoothing issue. This issue refers to the tendency of a model to produce indistinguishable features for nodes in different classes as the model depth increases, which consequently degrades the performance of the learning framework in downstream tasks. Consequently, to further support the effectiveness of SHNN, we theoretically analyse its robustness against the oversmoothing issue. Specifically, we demonstrate the following Proposition:
Proposition 4.2.
Let =$\mathcal{H}=\{\mathcal{V},\mathcal{E},\mathbf{H}\}$ denote a hypergraph, $\mathcal{G}=\{\mathcal{V},\mathbf{W}_{H}\}$ be its clique expansion with edge weights computed by Eq. (6), $\mathbf{L}$ be the graph Laplacian matrix of $\mathcal{G}$ computed by a symmetrically normalised and self-loops added $\mathbf{W}_{H}$, $\mathbf{X}_{\mathcal{V}}$ represent the input node features, and $\alpha\in[0,1)$ be a hyperparameter. We define that $F(\mathbf{X}^{\prime})=\operatorname{tr}(\mathbf{X}^{\prime\top}\mathbf{L}\mathbf{X}^{\prime})+\frac{\alpha}{1-\alpha}\operatorname{tr}[(\mathbf{X}^{\prime}-\mathbf{X}_{\mathcal{V}})^{\top}(\mathbf{X}^{\prime}-\mathbf{X}_{\mathcal{V}})]$, and $F_{min}$ as the global minimal value of $F(\mathbf{X}^{\prime})$ for $\mathbf{X}\in\mathbb{R}^{n\times d}$. Assume that: 1) $\hat{\mathbf{X}}_{\mathcal{V}}=\mathbf{S}\mathbf{X}_{\mathcal{V}}$; 2) $\mathbf{S}$ is computed by Eq. (7); 3) $\alpha>0$; and 4) $L\to+\infty$. Then, $F_{min}=F(\hat{\mathbf{X}}_{\mathcal{V}})$.
Proof Sketch.
We conduct this proof in two steps. First, we determine the close form for the global minimal point $\mathbf{X}^{\star}$ of $F(\mathbf{X}^{\prime})$. Second, we show that as $L\to+\infty$, the output of an $L$-layer SHNN converges to $\mathbf{X}^{\star}$. We elaborate on this proof in Appendix. ∎
Table 3: Dataset statistics. CE homophily is the homophily score based on clique expansions of hypergraphs; details of this score can be found in Appendix. Moreover, NC denotes the hypergraph node classification task and HP represents the hyperedge prediction task.
Cora-CA DBLP-CA Citeseer Congress House Senate
# Nodes 2708 41302 3312 1718 1290 282
# Hyperedges 1072 22363 1079 83105 340 315
# Features 1433 1425 3703 100 100 100
# Classes 7 6 6 2 2 2
CE homophily 0.803 0.869 0.893 0.555 0.498 0.509
Node Definition Paper Paper Paper Congressperson Representative Representative
Hyperedge Definition Co-authorship Co-authorship Co-citation Legislative-Bills-Sponsorship Committee Legislative-Bills-Sponsorship
Task NC&HP NC NC & HP NC NC&HP NC
Notably, minimising the first term of $F(\mathbf{X}^{\prime})$ enhances feature similarity among connected nodes on $\mathcal{H}$, and minimising the second term of $F(\mathbf{X}^{\prime})$ encourages the generated features of each node to retain the distinct information from its input features. Therefore, node features that can minimise $F(\mathbf{X}^{\prime})$ capture two key properties: 1) neighbouring nodes on $\mathcal{H}$ have similar features; and 2) each node contains unique information reflecting its individual input characteristics. Consequently, based on Proposition 4.2, an extremely deep SHNN tends to assign similar features to connected nodes on a hypergraph while preserving the unique input information of each node.
According to the discussion above, even if the number of layers in SHNN is increased to extremely large, SHNN can still preserve the unique input information of each node, which enables it to be robust against the oversmoothing issue. This property further ensures the effectiveness of our SHNN. We conduct extensive experiments in Section 6 to support the analysis presented in this section.
5 Related Work
Efficient learning with hypergraphs. Recent efforts have risen to utilize pre-defined hypergraph structures to enhance node feature learning, thereby benefiting downstream tasks (Bick et al., 2023). The predominant models in this area are hypergraph neural networks (HNNs). Inspired by the success of graph neural networks (GNNs) (Wu et al., 2020), most existing works focus on developing HNNs with a message passing mechanism that enables information exchange between connected nodes (Antelmi et al., 2023). Similar to their graph counterparts (Wu et al., 2019; Frasca et al., 2020), HNNs suffer from low learning efficiency. There mainly exist two recent works for this limitation (Feng et al., 2024; Tang et al., 2024). Specifically, (Feng et al., 2024) distills an HNN named HGNN (Feng et al., 2019) into an MLP during the training process, while (Tang et al., 2024) integrates hypergraph structural information into an MLP via a loss function computed by the sum of the maximum node feature distance in each hyperedge. These previous works focus on reducing model inference complexity via a dedicated training process, suitable for scenarios where training and test data are clearly separated. However, many hypergraph-related models are applied in semi-supervised node classification settings that involve the simultaneous usage of training and test data (Feng et al., 2019; Bai et al., 2021; Huang & Yang, 2021; Chien et al., 2022; Wang et al., 2023a, b; Duta et al., 2023; Telyatnikov et al., 2023; Joachims et al., 1999; Chapelle et al., 2009). Hence, reducing the training complexity of hypergraph-related models remains an important yet underexplored question.
We address this challenge by leveraging hypergraph structures via a training-free model. This idea is mainly inspired by a previous work in the graph domain named SIGN (Frasca et al., 2020), which moves the integration of graph information from the training stage to the data preprocessing phase based on training-free graph filters. The main difficulty in implementing this idea for hypergraphs lies in the requirement for hypergraph filters to operate with an incidence tensor (Zhang et al., 2019; Qu et al., 2022), resulting in prohibitively high memory usage for practical applications. Specifically, for a hypergraph with $n$ nodes and a maximum hyperedge size of $k$, the corresponding incidence tensor has $n^{k}$ dimensions. Therefore, instead of using hypergraph filters, we further draw inspiration from SGC (Wu et al., 2019) to introduce a a novel training-free model for hypergraphs, the simplified hypergraph neural network (SHNN). Unlike the SGC approach, which simplifies only the vanilla GCN model (Kipf & Welling, 2017) to create a pure low-pass filter that exclusively smooths the features of connected nodes, our method simplifies four representative Hypergraph Neural Networks (HNNs). This results in a model that not only smooths the features of connected nodes but also preserves the unique input information of each node.
6 Experiments
6.1 Experiment Setup
Task description. We conduct experiments in two downstream tasks: hypergraph node classification and hyperlink prediction. In hypergraph node classification (Wang et al., 2023a, b; Duta et al., 2023), we are given the hypergraph structure $\mathbf{H}$, node features $\mathbf{X}_{\mathcal{V}}$, and a set of labelled nodes with ground truth labels $\mathbf{Y}_{lab}=\{\mathbf{y}_{v}\}_{v\in\mathcal{V}_{lab}}$, where $\mathbf{y}_{v_{i}}\in\{0,1\}^{c}$ be a one-hot label. Our objective is to classify the unlabeled nodes within the given hypergraph. In hyperlink prediction (Chen & Liu, 2023), we are given node features $\mathbf{X}_{\mathcal{V}}$, a set of observed real and fake hyperedges $\mathcal{E}_{ob}$, and a set of potential hyperedges $\mathcal{E}_{p}$. The task requires the model to distinguish between real and fake hyperedges in $\mathcal{E}_{p}$ using $\mathbf{X}_{\mathcal{V}}$ and $\mathcal{E}_{ob}$.
Dataset and baseline. For the hypergraph node classification task, we conduct experiments on six real-world hypergraphs: Cora-CA, DBLP-CA, Citeseer, Congress, House, and Senate. We use nine HNNs as baselines: HGNN (Feng et al., 2019), HCHA (Bai et al., 2021), HNHN (Dong et al., 2020), UniGCNII (Huang & Yang, 2021), AllDeepSets (Chien et al., 2022), AllSetTransformer (Chien et al., 2022), PhenomNN (Wang et al., 2023b), ED-HNN (Wang et al., 2023a), and Deep-HGNN (Chen et al., 2022). For the hyperlink prediction task, we compare our SHNN with AllSetTransformer, PhenomNN, ED-HNN, and Deep-HGNN on three real-world hypergraphs: Cora-CA, Citeseer, and House. The dataset statistics are summarized in Table 3, with detailed descriptions of datasets and baselines available in Appendix.
Implementation. For hypergraph node classification, following previous works (Wang et al., 2023a, b; Duta et al., 2023), we use a 50%/25%/25% train/validation/test data split and adapt the baseline classification accuracy from these studies [^2] Since there are no reported results for Deep-HGNN under the chosen data split, we use our own results.. Additionally, similar to these works, we implement the classifier based on MLP for our SHNN and report the classification accuracy from ten runs. For hyperlink prediction, existing works (Chen & Liu, 2023) primarily focus on the design of the prediction head, with no reported results for applying our baseline HNNs to this task; so we report all results conducted by ourselves. We employ the deep set function implemented by (Chien et al., 2022) as the prediction head for our method and baseline methods due to its simplicity. We use a 50%/25%/25% train/validation/test data split and ensure each split contains five times as many fake hyperedges as real hyperedges. Moreover, both SHNN and HNNs only use real hyperedges in training and validation sets to do message passing. The learning frameworks are evaluated by Area Under the ROC curve (AUC). All experiments were performed on an RTX 3090 using PyTorch, and our code with running commands will be at https://anonymous.4open.science/r/SHNN. More details about the model hyperparameters are provided in Appendix.
6.2 Comparison with Baselines
Table 4: Node classification accuracy (%) for HNN-based and SHNN-based learning frameworks on benchmarks. The state-of-the-art result on each dataset is highlighted in bold font. The second and third highest accuracies are marked with an underline for easy identification.
Cora-CA DBLP-CA Citeseer Congress House Senate Avg. Mean
HGNN 82.64 ± 1.65 91.03 ± 0.20 72.45 ± 1.16 91.26 ± 1.15 61.39 ± 2.96 48.59 ± 4.52 74.56
HCHA 82.55 ± 0.97 90.92 ± 0.22 72.42 ± 1.42 90.43 ± 1.20 61.36 ± 2.53 48.62 ± 4.41 74.38
HNHN 77.19 ± 1.49 86.78 ± 0.29 72.64 ± 1.57 53.35 ± 1.45 67.80 ± 2.59 50.93 ± 6.33 68.12
UniGCNII 83.60 ± 1.14 91.69 ± 0.19 73.05 ± 2.21 94.81 ± 0.81 67.25 ± 2.57 49.30 ± 4.25 76.62
AllDeepSets 81.97 ± 1.50 91.27 ± 0.27 70.83 ± 1.63 91.80 ± 1.53 67.82 ± 2.40 48.17 ± 5.67 75.31
AllSetTransformer 83.63 ± 1.47 91.53 ± 0.23 73.08 ± 1.20 92.16 ± 1.05 51.83 ± 5.22 69.33 ± 2.20 76.93
PhenomNN 85.81 ± 0.90 91.91 ± 0.21 75.10 ± 1.59 88.24 ± 1.47 70.71 ± 2.35 67.70 ± 5.24 79.91
ED-HNN 83.97 ± 1.55 91.90 ± 0.19 73.70 ± 1.38 95.00 ± 0.99 72.45 ± 2.28 64.79 ± 5.14 80.30
Deep-HGNN 84.89 ± 0.88 91.76 ± 0.28 74.07 ± 1.64 93.91 ± 1.18 75.26 ± 1.76 68.39 ± 4.79 81.38
SHNN (ours) 86.54 ± 1.32 91.80 ± 0.30 74.82 ± 1.67 95.09 ± 0.89 76.29 ± 1.99 70.42 ± 2.74 82.50
Figure 2: Comparison of the training time required for HNN-based and SHNN-based learning frameworks to achieve optimal performance on benchmarks.
(a) Cora-CA.
(b) Citeseer.
(c) House.
Figure 3: Hyperlink prediction Area Under the ROC curve (AUC) (%) and relative training time for HNN-based and SHNN-based learning frameworks on benchmarks.
Node classification. We summarize the classification accuracy and relative training time of the SHNN-based and HNN-based frameworks in Table 1 and Figure 2, respectively. Table 4 demonstrates that SHNN not only leads to the best results across four datasets (Cora-CA, Congress, House, and Senate) but also results in the highest average mean accuracy overall. These findings confirm the effectiveness of SHNN in generating powerful node features for node classification. Further, as shown in Figure 2, the SHNN-based framework consistently needs less training time to achieve superior performance compared to the HNN-based framework. This underscores the efficacy of SHNN in enhancing training efficiency for learning frameworks in hypergraph-related downstream tasks.
Hyperlink prediction. The results for the hyperlink prediction task are summarized in Figure 3. These results exhibit a pattern similar to that observed in the hypergraph node classification task: the SHNN-based framework outperforms the HNN-based frameworks while requiring less training time. An interesting observation is that the training efficiency improvement brought by SHNN is smaller in the hyperlink prediction task compared to the node classification task. We attribute this to the fact that the task-specific model used for hyperlink prediction is more computationally expensive than the one used for hypergraph node classification. Specifically, the hyperlink prediction model needs to generate a prediction for each hyperedge by aggregating the features of the nodes, which has a complexity of $\mathcal{O}(nd_{v})$, whereas the node classification model only needs to perform forward propagation for each individual node, with a complexity of $\mathcal{O}(n)$. Here $n$ is the number of nodes and $d_{v}$ is the average node degree. Appendix theoretically demonstrates that the efficiency improvement provided by SHNN is inversely correlated with the complexity of the task-specific model.
6.3 Ablation study
We conduct a series of ablation studies on the SHNN-based learning framework in node classification on the Cora-CA, Citeseer, House, and Senate datasets.
(a) Cora-CA.
(b) Citeseer.
(c) House.
(d) Senate.
Figure 4: The ablation study on the number of layers and the $\alpha$ of SHNN in node classification.
Table 5: The ablation study on the SHNN-based node classification framework. The state-of-the-art result on each dataset is highlighted in bold font.
SHNN Weighted $\mathbf{S}$ Cora-CA Citeseer Senate House Avg. Mean
$\bm{\times}$$\bm{\times}$ 77.28 ± 1.57 73.89 ± 1.75 68.31 ± 3.40 74.43 ± 2.44 73.48
$\bm{\surd}$$\bm{\times}$ 84.90 ± 1.55 74.75 ± 1.62 68.87 ± 4.57 75.57 ± 1.81 76.02
$\bm{\surd}$ $\bm{\surd}$ 86.54 ± 1.32 74.82 ± 1.67 70.42 ± 2.74 76.29 ± 1.99 77.02
Ablation study on hyperparameters. We focus on two key hyperparameters of SHNN: the number of layers and the $\alpha$. As illustrated in Figure 4, models with $\alpha>0$ exhibit more consistent accuracy as the number of layers increases compared to models with $\alpha=0$. We attribute this consistency to the robustness discussed in Proposition 4.2, where a positive $\alpha$ helps retain the distinct information of each node, thereby mitigating the oversmoothing issue. Additionally, we observe that a positive $\alpha$ significantly enhances performance in the House and Senate datasets, which are more heterophilic, compared to the Cora-CA and Citeseer datasets. This suggests that retaining the distinct information of each node is more beneficial for heterophilic datasets.
Ablation study on learning framework components. In this study, we examine the influence of SHNN and hyperedge-size-based clique expansion on the performance of an SHNN-based node classification framework. Without SHNN, the framework functions as a pure MLP. We define $\mathbf{S}$ based on hyperedge-size-based clique expansion as weighted $\mathbf{S}$; in contrast, for the framework without weighted $\mathbf{S}$, $\mathbf{S}$ is generated using a $\mathbf{W}$ with all positive values set to one. As shown in Table 5, the inclusion of SHNN improves classification accuracy across all datasets, demonstrating that SHNN effectively incorporates hypergraph structural information into node features. Furthermore, adding weighted $\mathbf{S}$ further enhances performance, highlighting the informativeness of the hyperedge-size-based clique expansion. Thus, the components we designed significantly aid the task-specific model in leveraging useful information from hypergraph data.
7 Conclusion & Limitation
In this paper, we propose the Simplified Hypergraph Neural Network (SHNN), a novel training-free model designed specifically for hypergraph-structured data. SHNN addresses the challenge of balancing efficiency and effectiveness in the design of hypergraph machine learning models. We present both theoretical and empirical evidence demonstrating that SHNN can solve hypergraph-related downstream tasks with remarkable efficiency and effectiveness. Our SHNN can advance the field of hypergraph machine learning research by largely improving the efficiency of using observed hypergraph structures to solve downstream tasks.
Limitation and future work. Learning tasks on hypergraphs can generally be defined at the node, hyperedge, and hypergraph levels. In this paper, we empirically test our SHNN on node-level and hyperedge-level tasks, with the main limitation being the lack of empirical results on hypergraph-level tasks. To our knowledge, designing task-specific models to generate hypergraph-level features based on existing node features remains a challenging topic in the literature. Therefore, we consider the application of our SHNN to hypergraph-level tasks as future work.

### v3

@@
 Abstract
-Hypergraphs, with hyperedges connecting multiple nodes, are crucial for modelling higher-order interactions in real-world data.
-In frameworks utilising hypergraphs for downstream tasks, a task-specific model is typically paired with a hypergraph neural network (HNN).
-HNNs enhance the task-specific model by generating node features with hypergraph structural information via message passing.
-However, the training for HNNs is often computationally intensive, which limits their practical use.
-To tackle this challenge, we propose an alternative approach by integrating hypergraph structural information into node features using a training-free model called simplified hypergraph neural network (SHNN) that only contains a predefined propagation step.
-We theoretically show the efficiency and effectiveness of SHNN by showing that: 1) It largely reduces the training complexity when solving hypergraph-related downstream tasks compared to existing HNNs; 2) It utilises as much information as existing HNNs for node feature generation; and 3) It is robust against the oversmoothing issue while using long-range interactions.
-Experiments in node classification and hyperedge prediction showcase that, compared to state-of-the-art HNNs, SHNN leads to both competitive performance and superior training efficiency.
-Notably, on Cora-CA, the SHNN-based framework achieves the highest node classification accuracy with just $2\%$ training time of the best baseline.
+Hypergraphs are crucial for modeling higher-order interactions in real-world data.
+Hypergraph neural networks (HNNs) effectively utilise these structures by message passing to generate informative node features for various downstream tasks like node classification.
+However, the message passing block in existing HNNs typically requires a computationally intensive training process, which limits their practical use.
+To tackle this challenge, we propose an alternative approach by decoupling the usage of the hypergraph structural information from the model training stage.
+The proposed model, simplified hypergraph neural network (SHNN), contains a training-free message-passing block that can be precomputed before the training of SHNN, thereby reducing the computational burden.
+We theoretically support the efficiency and effectiveness of SHNN by showing that: 1) It is more training-efficient compared to existing HNNs; 2) It utilises as much information as existing HNNs for node feature generation; and 3) It is robust against the oversmoothing issue while using long-range interactions.
+Experiments based on six real-world hypergraph benchmarks in node classification and hyperlink prediction present that, compared to state-of-the-art HNNs, SHNN shows both competitive performance and superior training efficiency.
+Specifically, on Cora-CA, SHNN achieves the highest node classification accuracy with just $2\%$ training time of the best baseline.
 Machine Learning, ICML
@@
 Higher-order interactions involving more than two entities exist in various domains, such as co-authorships in social science (Han et al., 2009) and the spreading phenomena in epidemiology (Jhun, 2021).
-To model these complicated interactions, hypergraphs, an extension of traditional graphs, are widely employed (Bick et al., 2023).
-Hypergraphs consist of nodes representing entities and hyperedges denoting higher-order interactions that connect multiple entities.
-Recently, hypergraph neural networks (HNNs) have leveraged message passing over hypergraph structures to enhance node feature learning, achieving remarkable performance in various downstream tasks like node classification (Feng et al., 2019; Bai et al., 2021; Chien et al., 2022; Wang et al., 2023a; Huang & Yang, 2021; Wang et al., 2023b; Duta et al., 2023; Telyatnikov et al., 2023).
-However, the training of HNNs is typically computationally intensive, which limits their practical applications, especially in semi-supervised learning that utilises both training and test data simultaneously (Joachims et al., 1999; Chapelle et al., 2009).
-To address this challenge, we take inspiration from the literature on scalable and efficient graph neural networks (GNNs).
-In this literature, SIGN (Frasca et al., 2020) proposes to move the integration of graph information from the training stage to the data preprocessing phase by a training-free model, and SGC (Wu et al., 2019) builds an efficient architecture by collapsing multiple layers of graph convolutions with the intermediate non-linear activation functions removed.
-We aim to explore similar ideas for hypergraphs.
-In this paper, we propose a novel model called simplified hypergraph neural network (SHNN), by simplifying four representative HNNs in two steps.
-First, we make the selected HNNs training-free by eliminating their learnable parameters.
-Next, we enhance inference efficiency by removing non-linear activation functions and converting an $L$-layer HNN into a single propagation step.
-Remarkably, this two-step simplification unifies the four chosen HNNs into a single formulation, despite their different underlying design philosophies.
-We refer to this formulation as SHNN.
-We further conduct a thorough analysis of SHNN to theoretically demonstrate its efficiency and effectiveness.
-Generally, we present three key properties of SHNN: 1) SHNN significantly reduces training complexity in hypergraph-related downstream tasks compared to existing HNNs; See Figure 1 for the comparison between the training pipelines of an SHNN and an HNN in node classification; 2) SHNN can utilise as much information as existing HNNs for generating node features; and 3) SHNN is robust against the oversmoothness issue when taking into account the long-range informations.
+Hypergraphs, extending traditional graphs, effectively model these complex interactions with nodes representing entities and hyperedges connecting multiple entities (Bick et al., 2023).
+To leverage such structures for solving downstream tasks, the design of learning algorithms on hypergraphs has garnered increasing attention.
+Inspired by the success of graph neural networks (GNNs) (Wu et al., 2020), most current research focuses on developing hypergraph neural networks (HNNs) with a message passing block (MP-Block) and a task-specific block [^1] For instance, in node classification, the task-specific block is implemented as a classifier, whereas in hyperlink prediction, it is designed as a predictor.
+Following previous works (Feng et al., 2019; Bai et al., 2021; Chien et al., 2022; Wang et al., 2023a; Huang & Yang, 2021; Wang et al., 2023b; Duta et al., 2023; Telyatnikov et al., 2023), we focused on designing MP-Blocks that are compatible with various task-specific blocks..
+The MP-Block enables information exchange between connected nodes to generate informative node features for the task-specific block (Feng et al., 2019; Bai et al., 2021; Chien et al., 2022; Wang et al., 2023a; Huang & Yang, 2021; Wang et al., 2023b; Duta et al., 2023; Telyatnikov et al., 2023).
+However, similar to other message passing neural networks (Frasca et al., 2020; wu2022graph), training the MP-Block makes loss computation interdependent for connected nodes, resulting in a computationally intensive training process for existing HNNs.
+This limits their practical applications, especially in semi-supervised learning that utilises the training and test data simultaneously (Joachims et al., 1999; Chapelle et al., 2009).
+To address this challenge, inspired by the literature on efficient and effective GNNs (Frasca et al., 2020; Wu et al., 2019), our key solution is to move the integration of the hypergraph structural information from the training stage to the data preprocessing phase.
+This leads to a novel model called simplified hypergraph neural network (SHNN).
+In this model, a task-specific block is paired with a simplified message passing block (SMP-Block) that is training-free and can be precomputed in the data preprocessing phase.
+We develop the SMP-Block by simplifying the MP-Block of four state-of-the-art HNNs (Huang & Yang, 2021; Chen et al., 2022; Chien et al., 2022; Wang et al., 2023a) in two steps.
+First, we make the selected MP-Blocks training-free by eliminating their learnable parameters.
+Next, we enhance inference efficiency by removing non-linear activation functions and converting an $L$-layer MP-Blocks into a single propagation step.
+Remarkably, this two-step simplification unifies the chosen MP-Blocks into a single formulation, despite their different underlying design philosophies.
+We refer to this formulation as SMP-Block.
+Further, we thoroughly analyse SHNN to demonstrate its efficiency and effectiveness theoretically.
+Generally, we show three key properties: 1) SHNN is more training-efficient compared to existing HNNs; See Figure 1 for the comparison between the training pipelines of the HNN and our SHNN in node classification; 2) SHNN can utilise as much information as existing HNNs for generating node features; and 3) SHNN is robust against the oversmoothness issue when taking into account the long-range information.
 The contributions of this work are summarised as follows:
-$\bullet$ We develop SHNN, an efficient and effective training-free model for generating node features with the structural information in pre-defined hypergraphs.
-$\bullet$ We theoretically support the effectiveness and efficiency of SHNN by showing that: 1) It leads to remarkably low training complexity when solving hypergraph-related downstream tasks; 2) It utilises the same amount of information as existing HNNs for generating node features; and 3) It is robust against the oversmoothing issue when utilising long-range interactions.
+$\bullet$ We develop SHNN, an efficient and effective model for hypergraph-structured data.
+To our knowledge, SHNN is the first model to integrate the hypergraph structural information in the data preprocessing stage, significantly enhancing training efficiency.
+$\bullet$ We theoretically support the effectiveness and efficiency of SHNN by showing that: 1) It leads to remarkably low training complexity when solving hypergraph-related downstream tasks; 2) It utilises the same amount of information as existing HNNs for generating node features; and 3) It is robust against the oversmoothing issue when utilising long-range information.
 $\bullet$ We conduct extensive experiments in both node classification and hyperedge prediciton tasks to compare SHNN with nine state-of-the-art HNNs.
-The empirical results show that the SHNN-based framework exhibits both the competitive performance and superior training efficiency.
-Figure 1: Training pipeline of an HNN vs.
-an SHNN in node classification.
-Top row: The HNN uses a hypergraph structure to generate node features by $(L+1)$ learnable forward propagation blocks (FP Blocks), which are then used by a classifier, with the HNN and the classifier being trained simultaneously.
-Bottom row: The SHNN-based framework consists solely of the classifier being trained for node classification and the SHNN, which can be precomputed prior to classifier training.
+The empirical results show that the proposed SHNN exhibits both the competitive performance and superior training efficiency.
+Figure 1: Training pipeline of HNN vs.
+SHNN in node classification.
+Top row: HNN uses a hypergraph structure to generate node features by a learnable MP-Block, which are then used by a classifier, with the MP-Block and the classifier being trained together.
+For brevity, we omit the MLP in HNN for the input node features.
+Bottom row: SHNN consists solely of the classifier being trained for node classification and the SMP-Block, which can be precomputed prior to classifier training.
 2 Notation
-Hypergraphs.
+Hypergraph.
 Let =$\mathcal{H}=\{\mathcal{V},\mathcal{E},\mathbf{H}\}$ be a hypergraph, where $\mathcal{V}=\{v_{1},v_{2},\cdots,v_{n}\}$ is the node set, =$\mathcal{E}=\{e_{1},e_{2},\cdots,e_{m}\}$ is the hyperedge set, and $\mathbf{H}\in\{0,1\}^{n\times m}$ is an incidence matrix in which $\mathbf{H}_{ik}=1$ means that $e_{j}$ contains node $v_{i}$ and $\mathbf{H}_{ik}=0$ otherwise.
@@
 In the following discussions, we assume the hypergraph does not have isolated nodes or empty/duplicate hyperedges.
-Moreover, we define: 1) The distance between two nodes on a hypergraph is the number of hyperedges in the shortest path between them, e.g., the distance is one if they are directly connected by a hyperedge; 2) The $k$-hop neighbours of a node $v_{i}$ are all nodes with a distance of $k$ or less; and 3) The neighbouring nodes of a node are its 1-hop neighbours.
-Graphs.
+Moreover, we define: 1) The distance between two nodes on a hypergraph is the number of hyperedges in the shortest path between them, e.g., the distance is one if they are directly connected by a hyperedge; and 2) The $k$-hop neighbours of a node $v_{i}$ are all nodes with a distance of $k$ or less.
+Graph and clique expansion.
 Let $\mathcal{G}=\{\mathcal{V},\mathbf{W}\}$ be a graph, where $\mathcal{V}=\{v_{1},v_{2},\cdots,v_{n}\}$ is the node set, and $\mathbf{W}\in\mathbb{R}_{\geq 0}^{n\times n}$ is the adjacency matrix of $\mathcal{G}$ in which $\mathbf{W}_{ij}>0$ means that $v_{i}$ and $v_{j}$ are connected and $\mathbf{W}_{ij}=0$ otherwise.
@@
 Moreover, we denote the graph Laplacian of $\mathcal{G}$ as $\mathbf{L}=\mathbf{D}-\mathbf{W}$.
-Clique expansion.
 Given a hypergraph =$\mathcal{H}=\{\mathcal{V},\mathcal{E},\mathbf{H}\}$, its clique expansion is defined as a graph $\mathcal{G}=\{\mathcal{V},\mathbf{W}\}$, where $\mathcal{V}$ remains unchanged, and $\mathbf{W}_{ij}>0$ if and only if $v_{i}$ and $v_{j}$ are connected by a hyperedge on $\mathcal{H}$ and $\mathbf{W}_{ij}=0$ otherwise.
-Hence, each hyperedge in $\mathcal{H}$ is a clique in $\mathcal{G}$.
+Therefore, each hyperedge in $\mathcal{H}$ is a clique in $\mathcal{G}$.
 Other notations.
@@
 3 Methodology
-In this section, we first revisit existing hypergraph neural networks (HNNs), summarising these models along with their corresponding learning frameworks for downstream tasks into general formulations.
-Furthermore, we develop the simplified hypergraph neural network (SHNN) by removing the learnable parameters and the non-linear function from four representative HNNs.
-3.1 Revisiting Hypergraph Neural Networks
-Table 1: Overview for four state-of-the-art HNNs.
+In this section, we aim to develop an efficient and effective model for hypergraph-structured data.
+In Subsection 3.1, we present the overall design of our simplified hypergraph neural network.
+Subsection 3.2 elaborates on the development of the simplified message passing block.
+3.1 Overview
+General formulation.
+Inspired by a previous work named SIGN (Frasca et al., 2020) in the GNN literature, we design a model, named simplified hypergraph neural network (SHNN), that can integrate the information about hypergraph structures into the data preprocessing stage, thus also removing the reliance on it at training.
+Mathematically, the SHNN can be summarised as the following formulation:
+$\hat{\mathbf{Y}}=\varphi_{\mathbf{\Theta}}(\hat{\mathbf{X}}_{\mathcal{V}}),~{}~{}\hat{\mathbf{X}}_{\mathcal{V}}=S(\mathbf{X}_{\mathcal{V}},\mathbf{H}),$ (1)
+where $\varphi_{\mathbf{\Theta}}(\cdot)$ is a task-specific block, $\hat{\mathbf{Y}}$ denotes the task-specific output whose dimension is task-dependent [^2] For instance, in the node classification task, $\hat{\mathbf{Y}}\in\mathbb{R}^{n\times c}$ contains the logits for $c$ categories; and in the hyperlink prediction task, $\hat{\mathbf{Y}}\in\mathbb{R}^{m_{p}}$ contains the probability of $m_{p}$ potential hyperedges., $S(\cdot)$ denotes an MP-Block that contains only pre-defined parameters, allowing it to be precomputed before the training phase, and $\hat{\mathbf{X}}_{\mathcal{V}}\in\mathbb{R}^{n\times d}$ denote the features generated by $S(\cdot)$.
+In SIGN, $S(\cdot)$ is implemented by the combination of some graph filters.
+However, we cannot directly use existing hypergraph filters (Zhang et al., 2019; Qu et al., 2022) for two main reasons.
+Firstly, existing hypergraph filters are primarily designed for $k$-uniform hypergraphs, where all hyperedges have size $k$.
+Secondly, these filters require the use of an incidence tensor, which leads to prohibitively high memory usage for practical applications.
+Specifically, for a hypergraph with $n$ nodes and hyperedges of uniform size $k$, the corresponding incidence tensor has $n^{k}$ dimensions.
+Hence, instead of using existing methods, we draw inspiration from SGC (Wu et al., 2019) in the GNN literature to design a novel $S(\cdot)$ for hypergraphs, named simplified message passing block (SMP-Block), by simplifying MP-Blocks in state-of-the-art HNNs.
+3.2 Simplified Message Passing Block
+In this section, we first revisit existing hypergraph neural networks (HNNs), summarising these models and their MP-Blocks into general formulations.
+Furthermore, we develop the simplified message passing block (SMP-Block) by removing the learnable parameters and the non-linear function from the MP-Block in four state-of-the-art HNNs.
+3.2.1 Revisiting Hypergraph Neural Networks
+Table 1: Overview for four state-of-the-art HNNs and our SHNN.
 In this table, $\gamma_{U},\gamma_{E},\gamma_{D},\gamma^{\prime}_{l}\!\in\!(0,1)$ are hyperparameters and $\mathbf{I}\in\mathbb{R}^{d\times d}$ denotes an identity matrix.
-Name Type Message Passing Formulation
-UniGCNII (Huang & Yang, 2021) Direct Message Passing $\mathbf{X}_{\mathcal{V}}^{(l)}\!=\!\sigma\Bigg{(}\Big{(}(\!1\!-\!\gamma_{U}\!)\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}^{-1/2}\mathbf{H}\!\tilde{\mathbf{D}}_{\mathcal{H}^{\mathcal{E}}}^{-1/2}\mathbf{D}_{\mathcal{H}^{\mathcal{E}}}^{-1}\mathbf{H}^{\top}\mathbf{X}_{\mathcal{V}}^{(l-1)}\!+\!\gamma_{U}\mathbf{X}_{\mathcal{V}}^{(0)}\Big{)}\mathbf{\Theta}^{(l)}\Bigg{)}$
-Deep-HGNN (Chen et al., 2022) Direct Message Passing $\mathbf{X}_{\mathcal{V}}^{(l)}\!=\!\sigma\Bigg{(}\Big{(}(\!1\!-\!\gamma_{D}\!)\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}^{-1/2}\mathbf{H}\!\mathbf{D}_{\mathcal{H}^{\mathcal{E}}}^{-1}\mathbf{H}^{\top}\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}^{-1/2}\mathbf{X}_{\mathcal{V}}^{(l)}\!+\!\gamma_{D}\mathbf{X}_{\mathcal{V}}^{(0)}\Big{)}\Big{(}(1-\gamma^{\prime}_{D})\mathbf{I}+\gamma^{\prime}_{D}\mathbf{\Theta}^{(l)}\Big{)}\Bigg{)}$
-AllDeepSets (Chien et al., 2022) Indirect Message Passing $\mathbf{X}_{\mathcal{V}}^{(l)}=\operatorname{MLP}\Bigg{(}\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}^{-1}\mathbf{H}\operatorname{MLP}\Big{(}\mathbf{D}_{\mathcal{H}^{\mathcal{E}}}^{-1}\mathbf{H}^{\top}\operatorname{MLP}(\mathbf{X}_{\mathcal{V}}^{(l-1)})\Big{)}\Bigg{)}$
-ED-HNN (Wang et al., 2023a) Indirect Message Passing $\mathbf{X}_{\mathcal{V}}^{(l)}=\operatorname{MLP}\Bigg{[}(1-\gamma_{E})\operatorname{MLP}\Bigg{(}\mathbf{D}^{-1}_{\mathcal{H}^{\mathcal{V}}}\mathbf{H}\operatorname{MLP}\Big{(}\mathbf{D}^{-1}_{\mathcal{H}^{\mathcal{E}}}\mathbf{H}^{\top}\operatorname{MLP}(\mathbf{X}_{\mathcal{V}}^{(l-1)})\Big{)}\Bigg{)}+\gamma_{E}\mathbf{X}_{\mathcal{V}}^{(0)}\Bigg{]}$
-Model design.
+Moreover, $M$ is the training computational complexity of the task-specific block, $n$ is the node count, $m$ is the hyperedge count, $m^{\prime}$ is the number of edges in the clique expansion, $\|\mathbf{H}\|_{0}$ is the number of non-zero values in $\mathbf{H}$, $T$ is the number of training epochs, $L$ is the number of layers, and $d$ is the feature dimension.
+Name Type Hypergraph-Wise MP Function Training Computational Complexity
+UniGCNII (Huang & Yang, 2021) Direct MP $\mathbf{X}_{\mathcal{V}}^{(l)}\!=\!\sigma\Bigg{(}\Big{(}(\!1\!-\!\gamma_{U}\!)\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}^{-1/2}\mathbf{H}\!\tilde{\mathbf{D}}_{\mathcal{H}^{\mathcal{E}}}^{-1/2}\mathbf{D}_{\mathcal{H}^{\mathcal{E}}}^{-1}\mathbf{H}^{\top}\mathbf{X}_{\mathcal{V}}^{(l-1)}\!+\!\gamma_{U}\mathbf{X}_{\mathcal{V}}^{(0)}\Big{)}\mathbf{\Theta}^{(l)}\Bigg{)}$ $\mathcal{O}(M+TL(n+m+\|\mathbf{H}\|_{0})d+TLnd^{2})$
+Deep-HGNN (Chen et al., 2022) Direct MP $\mathbf{X}_{\mathcal{V}}^{(l)}\!=\!\sigma\Bigg{(}\Big{(}(\!1\!-\!\gamma_{D}\!)\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}^{-1/2}\mathbf{H}\!\mathbf{D}_{\mathcal{H}^{\mathcal{E}}}^{-1}\mathbf{H}^{\top}\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}^{-1/2}\mathbf{X}_{\mathcal{V}}^{(l)}\!+\!\gamma_{D}\mathbf{X}_{\mathcal{V}}^{(0)}\Big{)}\Big{(}(1-\gamma^{\prime}_{D})\mathbf{I}+\gamma^{\prime}_{D}\mathbf{\Theta}^{(l)}\Big{)}\Bigg{)}$ $\mathcal{O}(M+TLm^{\prime}d+TLnd^{2})$
+AllDeepSets (Chien et al., 2022) Indirect MP $\mathbf{X}_{\mathcal{V}}^{(l)}=\operatorname{MLP}\Bigg{(}\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}^{-1}\mathbf{H}\operatorname{MLP}\Big{(}\mathbf{D}_{\mathcal{H}^{\mathcal{E}}}^{-1}\mathbf{H}^{\top}\operatorname{MLP}(\mathbf{X}_{\mathcal{V}}^{(l-1)})\Big{)}\Bigg{)}$ $\mathcal{O}(M+TL\|\mathbf{H}\|_{0}d+TL(n+m)d^{2})$
+ED-HNN (Wang et al., 2023a) Indirect MP $\mathbf{X}_{\mathcal{V}}^{(l)}=\operatorname{MLP}\Bigg{[}(1-\gamma_{E})\operatorname{MLP}\Bigg{(}\mathbf{D}^{-1}_{\mathcal{H}^{\mathcal{V}}}\mathbf{H}\operatorname{MLP}\Big{(}\mathbf{D}^{-1}_{\mathcal{H}^{\mathcal{E}}}\mathbf{H}^{\top}\operatorname{MLP}(\mathbf{X}_{\mathcal{V}}^{(l-1)})\Big{)}\Bigg{)}+\gamma_{E}\mathbf{X}_{\mathcal{V}}^{(0)}\Bigg{]}$ $\mathcal{O}(M+TL\|\mathbf{H}\|_{0}d+TL(n+m)d^{2})$
+SHNN (Ours) Direct MP $\hat{\mathbf{X}}_{\mathcal{V}}=\mathbf{S}\mathbf{X}_{\mathcal{V}}$ $\mathcal{O}(M)$
+General formulation. An HNN for downstream tasks is formulated as the following equation:
+$\hat{\mathbf{Y}}=\varphi_{\mathbf{\Theta}}(\hat{\mathbf{X}}_{\mathcal{V}}),~{}~{}\hat{\mathbf{X}}_{\mathcal{V}}=\Phi_{\mathbf{\Theta}}(\mathbf{X}_{\mathcal{V}},\mathbf{H}),$ (2)
+where $\varphi_{\mathbf{\Theta}}(\cdot)$ is a task-specific block, $\hat{\mathbf{Y}}$ denotes the task-specific output whose dimension is task-dependent, $\Phi_{\mathbf{\Theta}}(\cdot)$ is a learnable message passing block, and $\hat{\mathbf{X}}_{\mathcal{V}}\in\mathbb{R}^{n\times d}$ denote the features generated by $\Phi_{\mathbf{\Theta}}(\cdot)$.
+The blocks $\varphi_{\mathbf{\Theta}}(\cdot)$ and $\Phi_{\mathbf{\Theta}}(\cdot)$ are trained together with the supervision given by the downstream tasks.
+We detail the message passing block in the next paragraph.
+Message passing block.
 Generally, existing HNNs are based on the message passing mechanism, adapted from the graph machine learning literature (Xu et al., 2019; Bronstein et al., 2021).
@@
 $\mathbf{x}_{v_{i}}^{(0)}=\operatorname{MLP}(\mathbf{x}_{v_{i}}),~{}~{}\mathbf{x}_{v_{i}}^{(l)}=\phi_{\mathbf{\Theta}}(\mathbf{x}_{v_{i}}^{(0)},\mathbf{x}_{v_{i}}^{(l-1)},\oplus_{v_{j}\in\mathcal{N}_{\mathcal{H}_{v_{i}}}}\mathbf{x}_{v_{j}}^{(l-1)}),$ (3)
-where $\phi_{\mathbf{\Theta}}(\cdot)$ denotes a learnable non-linear function that facilitates the message passing between neighbouring nodes, and $\mathcal{N}_{\mathcal{H}_{v_{i}}}$ is the set of $1$-hop neighbours of $v_{i}$ on the hypergraph.
+where $\phi_{\mathbf{\Theta}}(\cdot)$ denotes a learnable node-wise message passing function that facilitates the information exchange between neighbouring nodes, and $\mathcal{N}_{\mathcal{H}_{v_{i}}}$ is the set of neighbours of $v_{i}$ on the hypergraph.
 Specifically, $\phi_{\mathbf{\Theta}}(\cdot)$ can be categorized into two types: direct and indirect message passing functions.
@@
 Mathematically these approaches can be formulated as:
-$\mathbf{x}_{v_{i}}^{(l)}$ $=f_{3}\Big{(}f_{0}(\mathbf{x}_{v_{i}}^{(0)})+f_{1}(\mathbf{x}_{v_{i}}^{(l-1)})+\sum_{v_{j}\in\mathcal{N}_{\mathcal{G}_{v_{i}}}}f_{2}(\mathbf{x}^{(l-1)}_{v_{j}})\Big{)},$ (2a)
+$\mathbf{x}_{v_{i}}^{(l)}$ $=f_{3}\Big{(}f_{0}(\mathbf{x}_{v_{i}}^{(0)})+f_{1}(\mathbf{x}_{v_{i}}^{(l-1)})+\sum_{v_{j}\in\mathcal{N}_{\mathcal{H}_{v_{i}}}}f_{2}(\mathbf{x}^{(l-1)}_{v_{j}})\Big{)},$ (4a)
 $\mathbf{x}_{v_{i}}^{(l)}$ $=g_{2}\Big{(}g_{0}(\mathbf{x}_{v_{i}}^{(0)})+\mathbf{p}^{\top}g_{1}(\mathbf{x}_{v_{i}}^{(l-1)},\oplus_{v_{j}\in\mathcal{N}_{\mathcal{H}_{v_{i}}}}\mathbf{x}_{v_{j}}^{(l-1)})\Big{)},$ (4b)
-where $f_{0}(\cdot)$, $f_{1}(\cdot)$, $f_{2}(\cdot)$, $f_{3}(\cdot)$, $g_{0}(\cdot)$, $g_{1}(\cdot)$, $g_{2}(\cdot)$ represent seven learnable non-linear functions, and $\mathbf{p}\in\mathbb{R}^{n_{v}}$ is a learnable vector used to aggregate the features of virtual nodes generated by $g_{1}(\cdot)$.
+where $f_{0}(\cdot)$, $f_{1}(\cdot)$, $f_{2}(\cdot)$, $f_{3}(\cdot)$, $g_{0}(\cdot)$, $g_{1}(\cdot)$, $g_{2}(\cdot)$ represent seven learnable functions, and $\mathbf{p}\in\mathbb{R}^{n_{v}}$ is a vector used to aggregate the features of virtual nodes generated by $g_{1}(\cdot)$.
 Here, $n_{v}$ is a hyperparameter corresponding to the number of the virtual nodes, and Eq.
@@
 (4b) represent the direct and indirect approaches respectively.
-Further, we detail four representative HNNs in Table 1, and show how specific HNNs are special cases of Eq.
-(1) in Appendix.
-Application for downstream tasks. The pipeline of an HNN-based learning framework for downstream tasks can be formulated as follows:
-$\hat{\mathbf{Y}}=\varphi_{\mathbf{\Theta}}(\hat{\mathbf{X}}_{\mathcal{V}}),~{}~{}\hat{\mathbf{X}}_{\mathcal{V}}=\Phi_{\mathbf{\Theta}}(\mathbf{X}_{\mathcal{V}},\mathbf{H}),$ (3)
-where $\varphi_{\mathbf{\Theta}}(\cdot)$ is a task-specific model, $\hat{\mathbf{Y}}$ denotes the task-specific output whose dimension is task-dependent [^1] For instance, in the node classification task, $\hat{\mathbf{Y}}\in\mathbb{R}^{n\times c}$ contains the logits for $c$ categories; and in the hyperlink prediction task, $\hat{\mathbf{Y}}\in\mathbb{R}^{m_{p}}$ contains the probability of $m_{p}$ potential hyperedges., $\hat{\mathbf{X}}_{\mathcal{V}}\in\mathbb{R}^{n\times d}$ denote the features generated by an HNN, and $\Phi_{\mathbf{\Theta}}(\cdot)$ is an HNN that updates the node features with the forward propagation functions formulated as Eq.
-(1).
-The models $\varphi_{\mathbf{\Theta}}(\cdot)$ and $\Phi_{\mathbf{\Theta}}(\cdot)$ are trained together with the supervision given by the downstream tasks.
-As illustrated in Table 1, existing HNNs contain complex learnable message passing formulation, which makes the training of the HNN-based framework computationally expensive.
-3.2 Simplifying Hypergraph Neural Networks
-In this section, our goal is to introduce the simplified hypergraph neural network (SHNN) by simplifying the representative HNNs presented in Table 1.
+Further, we summarise four state-of-the-art HNNs in Table 1, and provide a more detailed review of existing HNNs in Appendix.
+In HNNs, an MP-Block typically consists of a combination of $L$ message-passing functions.
+In the following subsection, we introduce our SMP-Block by simplifying the MP-Blocks of HNNs in Table 1.
+3.2.2 Simplifying Hypergraph Neural Networks
+Our goal is to introduce the SMP-Block by simplifying the state-of-the-art HNNs presented in Table 1.
 Learnable parameters elimination.
-To develop a training-free model, we first eliminate the learnable parameters from the message passing formulation of four selected HNNs.
-Specifically, we replace the learnable matrices in their formulations with identity matrices.
-This simplification allows us to represent the simplified formulations of the chosen HNNs as follows:
+To develop a training-free MP-Block, we first eliminate the learnable parameters from the message passing functions of the selected HNNs.
+Specifically, we replace the learnable matrices in these functions with identity matrices.
+This simplification allows us to represent the simplified message passing functions as follows:
 $\mathbf{X}_{\mathcal{V}}^{(l)}$ $=\sigma\Big{(}(1-\gamma_{U})\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}^{-1/2}\mathbf{H}\!\tilde{\mathbf{D}}_{\mathcal{H}^{\mathcal{E}}}^{-1/2}\mathbf{D}_{\mathcal{H}^{\mathcal{E}}}^{-1}\mathbf{H}^{\top}\mathbf{X}_{\mathcal{V}}^{(l-1)}+\gamma_{U}\mathbf{X}_{\mathcal{V}}^{(0)}\Big{)},$ (5a)
@@
 Recently, some works in graph machine learning (Wu et al., 2019; Wang & Zhang, 2022) show that the message passing mechanism can operate more efficiently without non-linear activation functions while maintaining high effectiveness.
-Inspired by these works, we further remove the non-linearity from the formulations represented by Eqs.
-(4$\sim$4d), thereby transforming an L-layer HNN into a single propagation step.
-For the sake of brevity, we present the simplification results as Lemma 3.1 and provide the mathematical proof for this Lemma in Appendix.
-Lemma 3.1.
-Let $\alpha\in[0,1)$ be a hyperparameter, $\mathbf{X}_{\mathcal{V}}$ be the input node features, and $\hat{\mathbf{X}}_{\mathcal{V}}$ be the output of an $L$-layer UniGCNII/AllDeepSets/ED-HNN/Deep-HGNN.
-Assume that the learnable parameters and the non-linearity are removed from the model.
-Then, for any hypergraph =$\mathcal{H}=\{\mathcal{V},\mathcal{E},\mathbf{H}\}$, there exists a clique expansion $\mathcal{G}=\{\mathcal{V},\mathbf{W}\}$ to make $\hat{\mathbf{X}}_{\mathcal{V}}=\Big{(}(1-\alpha)^{L}\mathbf{W}^{L}+\alpha\sum_{l=0}^{L-1}(1-\alpha)^{l}\mathbf{W}^{l}\Big{)}\mathbf{X}_{\mathcal{V}}$.
-According to Lemma 3.1, we design the formulation of our SHNN as the following equation:
-$\hat{\mathbf{X}}=\mathbf{S}\mathbf{X}_{\mathcal{V}},$ (5)
-where $\mathbf{S}=(1-\alpha)^{L}\mathbf{W}^{L}+\alpha\sum_{l=0}^{L-1}(1-\alpha)^{l}\mathbf{W}^{l}$ and we define that the $L$ used to compute $\mathbf{S}$ is the number of layers of an SHNN.
-To intuitively understand the behaviour of an $L$-layer SHNN, we present the Lemma 3.2.
+Inspired by these works, we further remove the non-linearity from the functions represented by Eqs.
+(5$\sim$5d), thereby transforming an MP-Block with $L$ message passing functions into a single propagation step.
+For the sake of brevity, we present the simplification results as Proposition 3.1 and provide the mathematical proof for this Proposition in Appendix.
+Proposition 3.1.
+Let $\mathbf{X}_{\mathcal{V}}$ be input node features, $\hat{\mathbf{X}}_{\mathcal{V}}$ be the output of the MP-Block of an UniGCNII/AllDeepSets/ED-HNN/Deep-HGNN with $L$ MP functions, and $\alpha\!\in\![\!0,1\!)$.
+Assume that learnable parameters and the non-linearity are removed from the block.
+For any hypergraph =$\mathcal{H}\!=\!\!\{\!\mathcal{V},\!\mathcal{E},\!\mathbf{H}\!\}$, there exists a clique expansion $\mathcal{G}\!=\!\{\!\mathcal{V},\mathbf{W}\!\}$ to make $\hat{\mathbf{X}}_{\mathcal{V}}\!=\!\Big{(}(1\!-\!\alpha)^{L}\!\mathbf{W}^{L}\!+\!\alpha\!\sum_{l=0}^{L-1}(1\!-\!\alpha)^{l}\!\mathbf{W}^{l}\Big{)}\mathbf{X}_{\mathcal{V}}$.
+According to Proposition 3.1, we design the formulation of our SMP-Block as the following equation:
+$\hat{\mathbf{X}}=\mathbf{S}\mathbf{X}_{\mathcal{V}},$
+where $\mathbf{S}=(1-\alpha)^{L}\mathbf{W}^{L}+\alpha\sum_{l=0}^{L-1}(1-\alpha)^{l}\mathbf{W}^{l}$.
+We define that the $L$ used to compute $\mathbf{S}$ is the number of layers of an SMP-Block.
+To intuitively understand the behaviour of an $L$-layer SMP-Block, we present the Lemma 3.1.
 For conciseness, we show the proof of this Lemma in Appendix.
@@
 Assume that $\alpha\in[0,1)$, and $\mathbf{S}\!=\!(1-\alpha)^{L}\mathbf{W}^{L}+\alpha\sum_{l=0}^{L-1}(1-\alpha)^{l}\mathbf{W}^{l}$.
-Then, we have $\mathbf{S}_{ij}>0$ if and only if $v_{i}$ is a $L$-hop neighbour of $v_{j}$ on $\mathcal{H}$, and $\mathbf{S}_{ij}=0$ otherwise.
-By Lemma 3.2, an $L$-layer SHNN enables the information exchange between any node $v_{i}$ and its $L$-hop neighbours on the hypergraph.
+Then, for $i\neq j$, we have $\mathbf{S}_{ij}>0$ if and only if $v_{i}$ is an $L$-hop neighbour of $v_{j}$ on $\mathcal{H}$, and $\mathbf{S}_{ij}=0$ otherwise.
+By Lemma 3.1, an $L$-layer SMP-Block enables the information exchange between any node $v_{i}$ and its $L$-hop neighbours on the hypergraph.
 In the next paragraph, we detail the specific designs of $\mathbf{S}$.
-Operator design.
-To apply the SHNN to hypergraph-related downstream tasks, the key is designing $\mathbf{W}$ to generate $\mathbf{S}$.
-We achieve this using a hyperedge-size-based edge weight that is defined as:
+Operator design. To use the SMP-Block in our SHNN, the key is designing $\mathbf{W}$ to generate $\mathbf{S}$. We achieve this using a hyperedge-size-based edge weight that is defined as:
 $\mathbf{W}_{{H}_{ij}}=\sum_{k=1}^{m}\frac{\delta(v_{i},v_{j},e_{k})}{\mathbf{D}_{\mathcal{H}^{\mathcal{E}}_{kk}}},$ (6)
-where $\delta(\cdot)$ is a function that returns 1 if $e_{k}$ connect $v_{i}$ and $v_{j}$ and returns 0 otherwise.
+where $\delta(\cdot)$ is a function that returns 1 if $e_{k}$ connects $v_{i}$ and $v_{j}$ and returns 0 otherwise.
 The main intuition behind the design of this edge weight is that, under certain conditions, the value of $\mathbf{W}_{H_{ij}}$ is positively correlated with the probability of nodes $v_{i}$ and $v_{j}$ having the same label.
@@
 4 Analysis
-In this section, we begin by showing the efficiency of the proposed SHNN through a comparison of the training complexity of the SHNN-based framework with the HNN-based framework in hypergraph-related downstream tasks.
+In this section, we begin by showing the efficiency of the proposed SHNN by comparing its training complexity with the HNNs.
 Additionally, we demonstrate the effectiveness of our SHNN by analyzing its information utilisation and its robustness against the oversmoothing issue.
-Training complexity in downstream tasks. According to Eq. (5), the SHNN-based framework for downstream tasks can be summarised as follows:
-$\hat{\mathbf{Y}}=\varphi_{\mathbf{\Theta}}(\mathbf{S}\mathbf{X}_{\mathcal{V}}).$ (8)
+Training complexity in downstream tasks.
 As illustrated by Eq.
-(8), the SHNN-based framework for downstream tasks requires training only the task-specific model.
+(1), the SHNN in downstream tasks requires training only the task-specific block.
 Additionally, since the SHNN relies on a pre-defined one-step propagation operator, its computations can be handled during data pre-processing.
-Consequently, the training complexity of the SHNN-based framework is determined solely by the task-specific model.
+Consequently, the training complexity of the SHNN-based framework is determined solely by the task-specific block.
 In contrast, on the basis of Eq.
-(3), the HNN-based framework involves training both the HNN and the task-specific model.
-See Figure 1 for an example that compares the training pipelines of the SHNN-based and HNN-based framework.
-The HNN includes complex learnable operations that facilitate message passing between connected nodes on a hypergraph, inherently making the training of the HNN-based framework more complex than the SHNN-based framework.
-To provide a precise comparison, we summarize the training complexity of the SHNN-based framework and the four frameworks based on different state-of-the-art HNNs in Table 2.
-This table shows that the training complexity of the SHNN-based framework is consistently lower than that of the HNN-based frameworks, theoretically confirming the training efficiency of the SHNN-based framework.
-Table 2: Training computational complexity of SHNN/HNN-based frameworks.
-Here, $M$ denote the training computational complexity of the task-specific model, $n$ is the number of nodes, $m$ is the number of hyperedges, $m^{\prime}$ is the number of edges in the clique expansion, $\|\mathbf{H}\|_{0}$ is the number of non-zero entities in $\mathbf{H}$, $L$ is the number of layers, and $d$ is the feature dimension.
-SHNN UniGCNII AllDeepSets Deep-HGNN ED-HNN
-$\mathcal{O}(M)$ $\mathcal{O}(M+TL(n+m+\|\mathbf{H}\|_{0})d+TLnd^{2})$ $\mathcal{O}(M+TL\|\mathbf{H}\|_{0}d+TL(n+m)d^{2})$ $\mathcal{O}(M+TLm^{\prime}d+TLnd^{2})$ $\mathcal{O}(M+TL\|\mathbf{H}\|_{0}d+TL(n+m)d^{2})$
+(2), the HNN involve training both the HNN and the task-specific block.
+The HNN includes complex learnable operations that facilitate message passing between connected nodes on a hypergraph, inherently making the training of the HNN-based framework more complex than the SHNN.
+See Figure 1 for an example that qualitatively compares the training pipelines of the SHNN and HNN in node classification.
+To provide a quantitative comparison, we summarize the training computational complexity of the SHNN and the four state-of-the-art HNNs in Table 1.
+This table shows that the training complexity of the SHNN is consistently lower than that of the HNN, theoretically confirming the training efficiency of the proposed SHNN.
 Information utilisation.
 The core strength of existing HNNs lies in their ability to utilize the information embedded within hypergraph structures to generate powerful node features for downstream tasks.
-Consequently, we analyse the effectiveness of our proposed SHNN by comparing its information utilization capabilities with those of existing HNNs, which is summarised in Proposition 4.1.
+Consequently, we analyse the effectiveness of our proposed SHNN by comparing its information utilization capabilities with those of existing HNNs and conclude that our methods utilise the same amount of information as exising HNNs without any loss, which is summarised in Proposition 4.1.
 Proposition 4.1.
-Let $H_{0}$ denote the entropy of the information that an HNN utilizes by $L$ message passing function $\phi_{\mathbf{\Theta}}(\cdot)$ introduced in Eq.
-(1) for generating node features.
-Similarly, let $H_{1}$ denote the entropy of the information that an $L$-layer SHNN uses for the same purpose.
-Then, $H_{0}=H_{1}$.
-Proof Sketch.
-To prove this proposition, we first let $H_{0}^{v_{i}}$ be the entropy of information used by an HNN with $L$ message passing layers in generating the features of node $v_{i}$, $H_{1}^{v_{i}}$ be the entropy of information used by an $L$-layer SHNN for the same purpose, and $H_{2}^{v_{i}}$ denote the entropy of information within node $v_{i}$ and its $L$-hop neighbours.
-Theron, we prove this proposition by showing that $H_{0}^{v_{i}}=H_{2}^{v_{i}}$ and $H_{1}^{v_{i}}=H_{2}^{v_{i}}$.
-The detailed proof is presented in Appendix.
-∎
+Let $H_{0}^{v^{L}_{i}}$ be the entropy of information used by an HNN with $L$ message passing functions in generating the features of a node $v_{i}$, $H_{1}^{v^{L}_{i}}$ be the entropy of information used by an SHNN with an $L$-layer SMP-Block for the same purpose, and $H_{2}^{v^{L}_{i}}$ denote the entropy of information within node $v_{i}$ and its $L$-hop neighbours on the hypergraph.
+Then, $H_{0}^{v^{L}_{i}}=H_{1}^{v^{L}_{i}}=H_{2}^{v^{L}_{i}}$
 This proposition theoretically demonstrates that SHNN is as effective as existing HNNs in utilizing information from pre-defined hypergraphs to generate node features.
@@
 Robustness against the oversmoothing issue.
-According to Lemma 3.2, SHNN can leverage global interactions within the given hypergraph to generate node features by increasing the number of model layers.
-However, as highlighted in a recent study (Chen et al., 2022), message-passing-based models on hypergraphs may suffer from the oversmoothing issue.
+According to Lemma 3.1, SHNN can leverage global interactions within the given hypergraph to generate node features by deepening the SMP-Block.
+However, as highlighted in a recent study (Chen et al., 2022), the message-passing-based models on hypergraphs may suffer from the oversmoothing issue.
 This issue refers to the tendency of a model to produce indistinguishable features for nodes in different classes as the model depth increases, which consequently degrades the performance of the learning framework in downstream tasks.
 Consequently, to further support the effectiveness of SHNN, we theoretically analyse its robustness against the oversmoothing issue.
-Specifically, we demonstrate the following Proposition:
+Inspired by previous works in the graph machine learning literature (Yang et al., 2021; zhang2020revisiting), we demonstrate the following Proposition:
 Proposition 4.2.
@@
 (7); 3) $\alpha>0$; and 4) $L\to+\infty$.
-Then, $F_{min}=F(\hat{\mathbf{X}}_{\mathcal{V}})$.
-Proof Sketch.
-We conduct this proof in two steps.
-First, we determine the close form for the global minimal point $\mathbf{X}^{\star}$ of $F(\mathbf{X}^{\prime})$.
-Second, we show that as $L\to+\infty$, the output of an $L$-layer SHNN converges to $\mathbf{X}^{\star}$.
-We elaborate on this proof in Appendix.
-∎
+Then, $F(\hat{\mathbf{X}}_{\mathcal{V}})=F_{min}$.
+Notably, minimising the first term of $F(\mathbf{X}^{\prime})$ enhances feature similarity among connected nodes on $\mathcal{H}$, and minimising the second term of $F(\mathbf{X}^{\prime})$ encourages the generated features of each node to retain the distinct information from its input features.
+Therefore, node features that can minimise $F(\mathbf{X}^{\prime})$ capture two key properties: 1) neighbouring nodes on $\mathcal{H}$ have similar features; and 2) each node contains unique information reflecting its individual input characteristics.
+Based on Proposition 4.2, a deep SMP-Block tends to assign similar features to connected nodes on a hypergraph while maintaining each node’s unique input information.
+Hence, even with an extremely large number of layers, the SMP-Block can preserve the unique input information of each node, making it robust against the oversmoothing issue.
+This property further ensures the effectiveness of our SHNN.
+In Section 6, we empirically support the analysis presented in this section.
+5 Related Work
+Efficient learning with hypergraphs.
+Recent efforts have risen to utilize existing hypergraphs to enhance node feature learning, thereby benefiting downstream tasks (Bick et al., 2023).
+The predominant models in this area are hypergraph neural networks (HNNs) with a message passing block that enables information exchange between connected nodes (Antelmi et al., 2023).
+Similar to their graph counterparts (Wu et al., 2019; Frasca et al., 2020), HNNs suffer from low learning efficiency.
+There are two recent works for this limitation (Feng et al., 2024; Tang et al., 2024).
+(Feng et al., 2024) distills an HNN named HGNN (Feng et al., 2019) into an MLP during training, while (Tang et al., 2024) integrates hypergraph structural information into an MLP via a loss function computed by the sum of the maximum node feature distance in each hyperedge.
+These works focus on reducing model inference complexity via a dedicated training process, suitable for scenarios where training and test data are clearly separated.
+However, many existing models are applied to semi-supervised hypergraph node classification settings that involve the simultaneous use of training and test data (Feng et al., 2019; Bai et al., 2021; Huang & Yang, 2021; Chien et al., 2022; Wang et al., 2023a, b; Duta et al., 2023; Telyatnikov et al., 2023; Joachims et al., 1999; Chapelle et al., 2009).
+Therefore, reducing the training complexity of hypergraph-based machine learning models remains an important yet underexplored challenge.
+We address this challenge with a novel model named simplified hypergraph neural network (SHNN), which is the first to incorporate hypergraph structural information during data preprocessing.
+This integration significantly improves training efficiency.
 Table 2: Dataset statistics.
@@
 Moreover, NC denotes the hypergraph node classification task and HP represents the hyperedge prediction task.
-Cora-CA DBLP-CA Citeseer Congress House Senate
-# Nodes 2708 41302 3312 1718 1290 282
-# Hyperedges 1072 22363 1079 83105 340 315
-# Features 1433 1425 3703 100 100 100
-# Classes 7 6 6 2 2 2
-CE homophily 0.803 0.869 0.893 0.555 0.498 0.509
-Node Definition Paper Paper Paper Congressperson Representative Representative
-Hyperedge Definition Co-authorship Co-authorship Co-citation Legislative-Bills-Sponsorship Committee Legislative-Bills-Sponsorship
-Task NC&HP NC NC & HP NC NC&HP NC
-Notably, minimising the first term of $F(\mathbf{X}^{\prime})$ enhances feature similarity among connected nodes on $\mathcal{H}$, and minimising the second term of $F(\mathbf{X}^{\prime})$ encourages the generated features of each node to retain the distinct information from its input features.
-Therefore, node features that can minimise $F(\mathbf{X}^{\prime})$ capture two key properties: 1) neighbouring nodes on $\mathcal{H}$ have similar features; and 2) each node contains unique information reflecting its individual input characteristics.
-Consequently, based on Proposition 4.2, an extremely deep SHNN tends to assign similar features to connected nodes on a hypergraph while preserving the unique input information of each node.
-According to the discussion above, even if the number of layers in SHNN is increased to extremely large, SHNN can still preserve the unique input information of each node, which enables it to be robust against the oversmoothing issue.
-This property further ensures the effectiveness of our SHNN.
-We conduct extensive experiments in Section 6 to support the analysis presented in this section.
-5 Related Work
-Efficient learning with hypergraphs.
-Recent efforts have risen to utilize pre-defined hypergraph structures to enhance node feature learning, thereby benefiting downstream tasks (Bick et al., 2023).
-The predominant models in this area are hypergraph neural networks (HNNs).
-Inspired by the success of graph neural networks (GNNs) (Wu et al., 2020), most existing works focus on developing HNNs with a message passing mechanism that enables information exchange between connected nodes (Antelmi et al., 2023).
-Similar to their graph counterparts (Wu et al., 2019; Frasca et al., 2020), HNNs suffer from low learning efficiency.
-There mainly exist two recent works for this limitation (Feng et al., 2024; Tang et al., 2024).
-Specifically, (Feng et al., 2024) distills an HNN named HGNN (Feng et al., 2019) into an MLP during the training process, while (Tang et al., 2024) integrates hypergraph structural information into an MLP via a loss function computed by the sum of the maximum node feature distance in each hyperedge.
-These previous works focus on reducing model inference complexity via a dedicated training process, suitable for scenarios where training and test data are clearly separated.
-However, many hypergraph-related models are applied in semi-supervised node classification settings that involve the simultaneous usage of training and test data (Feng et al., 2019; Bai et al., 2021; Huang & Yang, 2021; Chien et al., 2022; Wang et al., 2023a, b; Duta et al., 2023; Telyatnikov et al., 2023; Joachims et al., 1999; Chapelle et al., 2009).
-Hence, reducing the training complexity of hypergraph-related models remains an important yet underexplored question.
-We address this challenge by leveraging hypergraph structures via a training-free model.
-This idea is mainly inspired by a previous work in the graph domain named SIGN (Frasca et al., 2020), which moves the integration of graph information from the training stage to the data preprocessing phase based on training-free graph filters.
-The main difficulty in implementing this idea for hypergraphs lies in the requirement for hypergraph filters to operate with an incidence tensor (Zhang et al., 2019; Qu et al., 2022), resulting in prohibitively high memory usage for practical applications.
-Specifically, for a hypergraph with $n$ nodes and a maximum hyperedge size of $k$, the corresponding incidence tensor has $n^{k}$ dimensions.
-Therefore, instead of using hypergraph filters, we further draw inspiration from SGC (Wu et al., 2019) to introduce a a novel training-free model for hypergraphs, the simplified hypergraph neural network (SHNN).
-Unlike the SGC approach, which simplifies only the vanilla GCN model (Kipf & Welling, 2017) to create a pure low-pass filter that exclusively smooths the features of connected nodes, our method simplifies four representative Hypergraph Neural Networks (HNNs).
-This results in a model that not only smooths the features of connected nodes but also preserves the unique input information of each node.
+Name # Nodes # Hyperedges # Features # Classes CE homophily Node Definition Hyperedge Definition Task
+Cora-CA 2708 1072 1433 7 0.803 Paper Co-authorship NC&HP
+DBLP-CA 41302 22363 1425 6 0.869 Paper Co-authorship NC
+Citeseer 3312 1079 3703 6 0.893 Paper Co-citation NC & HP
+Congress 1718 83105 100 2 0.555 Congressperson Legislative-Bills-Sponsorship NC
+House 1290 340 100 2 0.498 state-of-the-art Committee NC&HP
+Senate 282 315 100 2 0.509 state-of-the-art Legislative-Bills-Sponsorship NC
 6 Experiments
@@
 For the hyperlink prediction task, we compare our SHNN with AllSetTransformer, PhenomNN, ED-HNN, and Deep-HGNN on three real-world hypergraphs: Cora-CA, Citeseer, and House.
-The dataset statistics are summarized in Table 3, with detailed descriptions of datasets and baselines available in Appendix.
+The dataset statistics are summarized in Table 2, with detailed descriptions of datasets available in Appendix.
+Metric.
+Following previous works (Wang et al., 2023a; Chen & Liu, 2023), we evaluate models for node classification and hyperlink prediction with classification accuracy and the Area Under the ROC Curve (AUC), respectively.
+We evaluate the efficiency of the models by the relative training time $r_{f}$, defined as $r_{f}=t_{f}/t_{S}$, where $t_{f}$ is the training time to achieve optimal performance for the evaluated model, and $t_{s}$ is the training time for optimal performance of our SHNN.
 Implementation.
-For hypergraph node classification, following previous works (Wang et al., 2023a, b; Duta et al., 2023), we use a 50%/25%/25% train/validation/test data split and adapt the baseline classification accuracy from these studies [^2] Since there are no reported results for Deep-HGNN under the chosen data split, we use our own results..
-Additionally, similar to these works, we implement the classifier based on MLP for our SHNN and report the classification accuracy from ten runs.
+For node classification, we follow previous works (Wang et al., 2023a; Duta et al., 2023) to use a 50%/25%/25% train/validation/test data split and adapt the baseline classification accuracy from them [^3] Since there are no reported results for Deep-HGNN under the chosen data split, we use our own results..
+Additionally, similar to these works, we implement the classifier based on MLP for our SHNN and report the results from ten runs.
 For hyperlink prediction, existing works (Chen & Liu, 2023) primarily focus on the design of the prediction head, with no reported results for applying our baseline HNNs to this task; so we report all results conducted by ourselves.
@@
 We use a 50%/25%/25% train/validation/test data split and ensure each split contains five times as many fake hyperedges as real hyperedges.
-Moreover, both SHNN and HNNs only use real hyperedges in training and validation sets to do message passing.
-The learning frameworks are evaluated by Area Under the ROC curve (AUC).
-All experiments were performed on an RTX 3090 using PyTorch, and our code with running commands will be at https://anonymous.4open.science/r/SHNN.
-More details about the model hyperparameters are provided in Appendix.
+Both SHNN and HNNs only use real hyperedges in training and validation sets to do message passing.
+All experiments were performed on an RTX 3090 with PyTorch, and our code with running commands is at https://anonymous.4open.science/r/SHNN.
+More details are in Appendix.
 6.2 Comparison with Baselines
-Table 4: Node classification accuracy (%) for HNN-based and SHNN-based learning frameworks on benchmarks.
-The state-of-the-art result on each dataset is highlighted in bold font.
-The second and third highest accuracies are marked with an underline for easy identification.
+Table 3: Node classification accuracy (%) for HNNs and SHNN. The best result on each dataset is highlighted in bold font. The second and third highest accuracies are marked with an underline.
 Cora-CA DBLP-CA Citeseer Congress House Senate Avg. Mean
@@
 SHNN (ours) 86.54 ± 1.32 91.80 ± 0.30 74.82 ± 1.67 95.09 ± 0.89 76.29 ± 1.99 70.42 ± 2.74 82.50
-Figure 2: Comparison of the training time required for HNN-based and SHNN-based learning frameworks to achieve optimal performance on benchmarks.
+Figure 2: The relative training time required for HNNs and SHNN to achieve optimal performance.
 (a) Cora-CA.
@@
 (c) House.
-Figure 3: Hyperlink prediction Area Under the ROC curve (AUC) (%) and relative training time for HNN-based and SHNN-based learning frameworks on benchmarks.
+Figure 3: Hyperlink prediction AUC (%) and relative training time for HNNs and SHNN.
 Node classification.
-We summarize the classification accuracy and relative training time of the SHNN-based and HNN-based frameworks in Table 1 and Figure 2, respectively.
+We summarize the classification accuracy and relative training time of the SHNN and HNNs in Table 3 and Figure 2, respectively.
 Table 3 demonstrates that SHNN not only leads to the best results across four datasets (Cora-CA, Congress, House, and Senate) but also results in the highest average mean accuracy overall.
 These findings confirm the effectiveness of SHNN in generating powerful node features for node classification.
-Further, as shown in Figure 2, the SHNN-based framework consistently needs less training time to achieve superior performance compared to the HNN-based framework.
-This underscores the efficacy of SHNN in enhancing training efficiency for learning frameworks in hypergraph-related downstream tasks.
+As discussed in Section 4, we attribute this effectiveness to SHNN’s ability to match the power of existing HNNs in utilizing neighbourhood information to generate node features, while requiring the optimization of fewer parameters.
+This makes the SHNN more efficient in utilizing training data and reduces the risk of overfitting, thereby outperforming more complex HNN counterparts.
+On the other hand, as shown in Figure 2, the SHNN consistently needs less training time to achieve superior performance compared to the HNNs.
+This underscores the high training efficiency of SHNN in downstream tasks.
 Hyperlink prediction.
 The results for the hyperlink prediction task are summarized in Figure 3.
-These results exhibit a pattern similar to that observed in the hypergraph node classification task: the SHNN-based framework outperforms the HNN-based frameworks while requiring less training time.
+These results exhibit a pattern similar to that observed in the hypergraph node classification task: the SHNN framework outperforms the HNNs while requiring less training time.
 An interesting observation is that the training efficiency improvement brought by SHNN is smaller in the hyperlink prediction task compared to the node classification task.
-We attribute this to the fact that the task-specific model used for hyperlink prediction is more computationally expensive than the one used for hypergraph node classification.
-Specifically, the hyperlink prediction model needs to generate a prediction for each hyperedge by aggregating the features of the nodes, which has a complexity of $\mathcal{O}(nd_{v})$, whereas the node classification model only needs to perform forward propagation for each individual node, with a complexity of $\mathcal{O}(n)$.
-Here $n$ is the number of nodes and $d_{v}$ is the average node degree.
-Appendix theoretically demonstrates that the efficiency improvement provided by SHNN is inversely correlated with the complexity of the task-specific model.
+The intuition behind this observation is that the task-specific block used for hyperlink prediction is more computationally expensive than the one used for hypergraph node classification.
+Specifically, the hyperlink predictor needs to generate a prediction for each hyperedge by aggregating the features of the nodes connected by the hyperedge, whereas the node classifier only needs to perform forward propagation for each individual node.
+We theoretically discuss that the efficiency improvement provided by SHNN is inversely correlated with the complexity of the task-specific block in Appendix.
 6.3 Ablation study
-We conduct a series of ablation studies on the SHNN-based learning framework in node classification on the Cora-CA, Citeseer, House, and Senate datasets.
+We conduct a series of ablation studies on both the SPM-Block and the general SHNN model in node classification on the Cora-CA, Citeseer, House, and Senate datasets.
 (a) Cora-CA.
@@
 (d) Senate.
-Figure 4: The ablation study on the number of layers and the $\alpha$ of SHNN in node classification.
-Table 5: The ablation study on the SHNN-based node classification framework. The state-of-the-art result on each dataset is highlighted in bold font.
-SHNN Weighted $\mathbf{S}$ Cora-CA Citeseer Senate House Avg. Mean
+Figure 4: The ablation study on the hyperparameters of SPM-Block in node classification.
+Table 4: The ablation study on the components of SHNN in node classification. The best result on each dataset is highlighted in bold font.
+SMP-Block Weighted $\mathbf{S}$ Cora-CA Citeseer Senate House Avg. Mean
 $\bm{\times}$$\bm{\times}$ 77.28 ± 1.57 73.89 ± 1.75 68.31 ± 3.40 74.43 ± 2.44 73.48
@@
 $\bm{\surd}$ $\bm{\surd}$ 86.54 ± 1.32 74.82 ± 1.67 70.42 ± 2.74 76.29 ± 1.99 77.02
-Ablation study on hyperparameters.
-We focus on two key hyperparameters of SHNN: the number of layers and the $\alpha$.
+Ablation study on the hyperparameters of SPM-Block.
+We focus on two key hyperparameters of the SPM-Block: the number of layers and the $\alpha$.
 As illustrated in Figure 4, models with $\alpha>0$ exhibit more consistent accuracy as the number of layers increases compared to models with $\alpha=0$.
@@
 This suggests that retaining the distinct information of each node is more beneficial for heterophilic datasets.
-Ablation study on learning framework components.
-In this study, we examine the influence of SHNN and hyperedge-size-based clique expansion on the performance of an SHNN-based node classification framework.
-Without SHNN, the framework functions as a pure MLP.
+Ablation study on the components of SHNN.
+In this study, we examine the influence of the SMP-Block and the hyperedge-size-based clique expansion on the performance of the SHNN in node classification.
+Without the SMP-Block, the model functions as a pure MLP.
 We define $\mathbf{S}$ based on hyperedge-size-based clique expansion as weighted $\mathbf{S}$; in contrast, for the framework without weighted $\mathbf{S}$, $\mathbf{S}$ is generated using a $\mathbf{W}$ with all positive values set to one.
-As shown in Table 5, the inclusion of SHNN improves classification accuracy across all datasets, demonstrating that SHNN effectively incorporates hypergraph structural information into node features.
+As shown in Table 4, the inclusion of SMP-Block improves classification accuracy across all datasets, demonstrating that SMP-Block effectively incorporates hypergraph structural information into node features.
 Furthermore, adding weighted $\mathbf{S}$ further enhances performance, highlighting the informativeness of the hyperedge-size-based clique expansion.
-Thus, the components we designed significantly aid the task-specific model in leveraging useful information from hypergraph data.
+Thus, the components we designed significantly aid SHNN in leveraging useful information from the given hypergraph structure.
 7 Conclusion & Limitation
-In this paper, we propose the Simplified Hypergraph Neural Network (SHNN), a novel training-free model designed specifically for hypergraph-structured data.
-SHNN addresses the challenge of balancing efficiency and effectiveness in the design of hypergraph machine learning models.
-We present both theoretical and empirical evidence demonstrating that SHNN can solve hypergraph-related downstream tasks with remarkable efficiency and effectiveness.
-Our SHNN can advance the field of hypergraph machine learning research by largely improving the efficiency of using observed hypergraph structures to solve downstream tasks.
+In this paper, we propose a novel model called the simplified hypergraph neural network (SHNN).
+The key innovation of SHNN is an original training-free messgea passing block tailored for data on hypergraphs.
+We present both theoretical and empirical results demonstrating that SHNN can solve hypergraph-related downstream tasks with remarkable efficiency and effectiveness.
+To our knowledge, SHNN is the first model to integrate the hypergraph structural information in the data preprocessing stage, which significantly enhances training efficiency.
+The proposed SHNN can advance the field of hypergraph machine learning research by not only improving the efficiency of using observed hypergraph structures to solve downstream tasks, but also serving as a simple starting point for the development of future hypergraph machine learning models.
 Limitation and future work.
@@
 In this paper, we empirically test our SHNN on node-level and hyperedge-level tasks, with the main limitation being the lack of empirical results on hypergraph-level tasks.
-To our knowledge, designing task-specific models to generate hypergraph-level features based on existing node features remains a challenging topic in the literature.
+To our knowledge, designing task-specific blocks to generate hypergraph-level features based on existing node features itself remains a challenging topic in the literature.
 Therefore, we consider the application of our SHNN to hypergraph-level tasks as future work.